### v3

# largesymbols"0D largesymbols"0D

Abstract
We explore pretraining estimators for structural econometric models. The estimator is “pretrained” in the sense that the bulk of the computational cost and researcher effort occur during its construction. Subsequent applications of the estimator to different datasets require little computational cost or researcher effort. The estimation leverages a neural net to recognize the structural model’s parameter from data patterns. As an initial trial, this paper builds a pretrained estimator for a sequential search model that is known to be difficult to estimate. We evaluate it on 12 real datasets. The estimation takes seconds to run and shows high accuracy. We provide the estimator at pnnehome.github.io. More generally, pretrained, off-the-shelf estimators can make structural models more accessible to researchers and practitioners.
1 Introduction
Pre-trained machine learning models, ranging from popular image recognition models to large language models, are achieving significant success. These pretrained models are particularly good at pattern-driven tasks that are otherwise laborious for humans. By leveraging pretrained models, end users can greatly reduce the costs of these tasks. Take image recognition for example: instead of collecting their own images, labeling them, and training a neural net from scratch, users can take a pretrained neural net and directly apply it to recognize objects in their images.
This paper explores bringing pretraining to structural estimation. We develop an approach to pretrain estimators for structural econometric models. The approach builds on Wei and Jiang (2024) (more discussion later). The estimator is “pretrained” in the sense that the bulk of computational cost and researcher effort occur during the construction of the estimator. Subsequent applications of it to different datasets require little cost or effort. A pretrained estimator assumes a given structural model. Then, it uses a neural net to “recognize” the structural model’s parameter from data patterns (similar to how neural nets recognize objects in images). Therefore, we refer to it as the pretrained neural net estimator (hereafter, pretrained NNE).
Pretrained NNE makes structural modeling more accessible. Estimating a structural model is often technically complex and time-consuming. Yet, the values of structural models – such as insights on micro-founded consumer preferences and counterfactual analyses – come only after the estimation. Pretrained NNE provides an off-the-shelf tool to easily estimate a structural model. It benefits researchers and practitioners who seek to apply a structural model without delving into the technical complexities of estimation. The ease of computation can also facilitate new applications, such as integrating structural models in real-time algorithms (e.g., dynamic pricing, recommenders, bandit experiments). Pretrained estimators can be called continuously to process real-time data and feed into algorithms to adjust prices, rankings, recommendations, and more.
As an initial trial, this paper focuses on sequential search models that are known to be difficult to estimate. These models predict which products a consumer will search and buy (e.g., searching products $j=1,2,8$ and buying $j=2$). One example of their applications is to model online shopping behaviors. Search models are commonly estimated by simulated maximum likelihood (SMLE), but the estimation is challenging. The vast number of search and purchase combinations makes direct likelihood simulation computationally infeasible. Workarounds have been proposed, such as likelihood smoothing or specialized simulation algorithms (see Ursu et al. 2023). Nevertheless, the computational cost of a single estimation run is still high. Moreover, multiple estimation runs are required to tune optimization routines, try starting points, tune smoothing factors, etc. Additional effort is needed to derive and code the complex simulation algorithm. More broadly, many other structural models face similar challenges in estimation.
Pretrained NNE offers an alternative way for researchers to estimate a structural model. It does so by learning a direct mapping from data patterns to the structural model’s parameter. Conceptually, we can think of a pretrained NNE as follows. Many papers of structural estimation start with reduced-form evidence before estimating the structural model. Although reduced-form results do not directly translate into an estimate of the structural model’s parameter, they offer useful clues. For example, if customers are more likely to search lower-price items, we expect the price coefficient in the search model to be more negative. If customers make more searches over weekends, we know weekend customers have either a low search cost or a low outside utility. More generally, we can imagine that with a rich set of such reduced-form patterns, an expert familiar with the search model can “guesstimate” the search model parameter value. The present paper’s exercise is to leverage machine learning to create such an “expert.”
Figure 1: An Illustration of NNE
We describe a standard sequential search model in Section 2. Then, in Section 3 we describe how to pretrain a NNE for this search model. At a high level, the training is based on a large number of datasets simulated from the search model. Let $\boldsymbol{\theta}$ denote the search model parameter. For any value of $\boldsymbol{\theta}$ drawn from a prior, we can use the search model to simulate a dataset ${\cal D}$. Use $\boldsymbol{m}$ to collect some data patterns of ${\cal D}$. Repeating this process $L$ times gives us $\{\boldsymbol{\theta}^{(\ell)},\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$. (We currently use $L\simeq$ 1 million). These $L$ pairs serve as the training examples for us to train a neural net mapping from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. This trained neural net is the aforementioned “expert” capable of estimating the search model. Figure 1 provides an illustration of the neural net.
We provide the pretrained NNE at: pnnehome.github.io. To apply it to estimate the search model, users only need to supply data. No further tuning is required, and the computation typically takes seconds on a laptop. Specifically, users supply a dataset ${\cal D}$ consisting of two components: (i) $\boldsymbol{y}$, which includes search and purchase choices, and (ii) $\boldsymbol{x}$, which includes product attributes that affect utility, consumer attributes that affect outside utility, and advertising attributes that affect search cost. The current version of our pretrained NNE can accommodate datasets with $n\geq 1000$ consumers, 2 to 8 product attributes, 0 to 5 consumer attributes, and 0 to 2 marketing attributes. These ranges are configurable before pretraining.
Section 4 evaluates the pretrained NNE using 12 real datasets. We conduct two sets of tests. The first set consists of Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices under a “true” $\boldsymbol{\theta}$. Pretrained NNE is faster than SMLE by 3-4 orders of magnitude. Further, it recovers $\boldsymbol{\theta}$ with accuracy equal to or higher than SMLE. The result holds even when compared against the SMLE that uses a specialized simulation algorithm for search models that takes lots of effort to implement (Jiang et al. 2021). The second set of tests estimates the search model entirely from real data. Because the true $\boldsymbol{\theta}$ is unknown, we compare the estimates by pretrained NNE with those by SMLE. Their estimates are similar. The pretrained NNE is again faster by 3-4 orders of magnitude.
Compared to the general practice of training image or language models, the pretraining of NNE exploits the specific context of structural estimation in two key aspects. First, image or language models typically must learn representations of visual or textual data to perform tasks. In contrast, we use economic theory and intuition to directly specify a data representation, $\boldsymbol{m}$, for estimating $\boldsymbol{\theta}$. For the search model, we let $\boldsymbol{m}$ include the coefficients from a set of regressions describing consumer searches and purchases. This greatly reduces the training complexity. Second, image or language models are trained with labeled images or written texts. In contrast, we generate training examples with the structural model. Specifically, we use the search model to simulate outcomes $\boldsymbol{y}$ given attributes $\boldsymbol{x}$, and we simulate $\boldsymbol{x}$ to reflect the attributes commonly seen in real applications. In particular, we use a broad range of specifications for $\boldsymbol{x}$ that vary in marginal distributions, correlations, and the sizes of within- and cross-consumer variations.
This paper extends the standard (non-pretrained) NNE in Wei and Jiang (2024). In the standard NNE, the training examples assume the observed attributes $\boldsymbol{x}$ from a given real dataset. Thus, the trained NNE is applicable specifically to that real dataset. A different real dataset requires training a separate NNE. In contrast, a pretrained NNE can be directly applied to different real datasets. [^1] Another extension to Wei and Jiang (2024) is that the current paper expands $\boldsymbol{m}$ to include not only data moments but also coefficients of reduced-form regressions. The expansion substantially improves estimation accuracy. Here, it is important to note that like the standard NNE, a pretrained NNE still assumes a specific structural model. Because the cost of building a pretrained NNE is considerably higher than training a standard NNE, pretrained estimators are perhaps best suited for structural models that are already well established in literature.
From a theoretical perspective, the estimate of NNE reflects a Bayesian posterior. Building on the results in Wei and Jiang (2024), one can show that as $L\rightarrow\infty$, the pretrained NNE gives the posterior of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. This posterior is one’s best estimate of the structural model parameter after seeing (only) the data patterns. It is also known as a “limited-information” posterior, because it conditions on a set of data patterns instead of full data.
Finally, we emphasize that the intention of this paper is to explore the feasibility of pretrained estimators for structural models, rather than to develop an all-around tool for estimating search models. So for the latter task, there remains room for improvement. For example, how to incorporate richer unobserved heterogeneity across consumers? How to allow instrumental variables to address possible endogeneity? Section 5 discusses these potential extensions.
1.1 Literature
This paper is related to several streams of research. First, this paper joins a recent literature that seeks to leverage machine learning for econometric estimation. In particular, the literature has studied the use of flexible machine learning models to better capture individual-level heterogeneity (e.g., Wager and Athey 2018, Athey et al. 2019, Farrell et al. 2021b), to better incorporate high-dimensional nuisance parameters (e.g., Chernozhukov et al. 2018), and to construct instrumental variables (e.g., Hartford et al. 2017, Lewis and Syrgkanis 2018, Singh et al. 2020). A main theme of this literature so far is to use machine learning to flexibly describe the relations between variables in data. The current paper differs in that the relations between data variables are described by an existing econometric model (e.g., sequential search), and we leverage machine learning to estimate the parameter of this econometric model. Recent works sharing this theme include Wei and Jiang (2024) and Kaji et al. (2023). In particular, Kaji et al. (2023) takes a substantially different approach. They borrow ideas from adversarial learning, and propose training a discriminator to construct the objective function for estimating the econometric model.
Second, we build on the successful idea of pretraining in machine learning practice. Pretraining involves training a machine learning model to learn general capabilities which can then be applied to various specific tasks. In natural language processing, pretrained models like BERT and GPT demonstrate remarkable capabilities in many language tasks (Devlin et al. 2018, Brown et al. 2020). In computer vision, pretrained neural nets such as AlexNet and ResNet have been applied widely (Krizhevsky et al. 2012, He et al. 2016). A review can be found in Han et al. (2021). Overall, pretrained models have considerably improved the productivity of researchers and practitioners. The current paper aims to bring the benefits of pretraining to structural estimation in economics and marketing. Pretrained estimators are “Train Once, Deploy Anywhere.” Their availability will make structural estimation more accessible to researchers and practitioners.
Third, this paper adds to the literature that seeks to enable model estimation while preserving data privacy. Data privacy becomes an increasingly important issue to marketers. Various methods have been proposed to preserve privacy in estimation, such as noise infusion (Abowd et al. 2012, Avella-Medina 2021), generating synthetic data (Schneider et al. 2018, Anand and Lee 2023), and estimation using aggregate data (e.g., Musalem et al. 2009). The present paper advances privacy-preserving estimation using aggregate data. Pretrained NNE uses only reduced-form data patterns to estimate a structural model. Sharing reduced-form data patterns, which are aggregated across individuals, poses far less privacy concern for data owners than sharing individual-level data.
2 Consumer search model
This section describes a standard (sequential) search model, which will be the focus of this paper. We also describe the typical data used to estimate the search model.
2.1 The model
Sequential search models describe how consumers search and choose from a set of available options. One example of the context where these models are increasingly applied is online browsing and shopping. In this context, consumer is given a webpage with a list of products; she decides which products to search (i.e., click) and which product to buy. A survey of search models can be found in Ursu et al. (2023).
The current paper focuses on a fairly standard search model. Formally, there are $n$ consumers (or search sessions) and $J$ products per consumer. Let vector $\boldsymbol{x}_{ij}^{(prod)}$ collect the attributes of product $j$ for consumer $i$. The utility of consumer $i$ from product $j$ is:
$u_{ij}=\boldsymbol{\beta}^{\prime}\boldsymbol{x}_{ij}^{(prod)}+\mu_{ij}+\varepsilon_{ij}.$
There are two shocks, $\mu_{ij}$ and $\varepsilon_{ij}$. The consumer observes $\mu_{ij}$ pre-search but observes $\varepsilon_{ij}$ only post-search. Both shocks are distributed as $N(0,1)$. While normalizing one shock is without loss of generality, normalizing both shocks is not. Nevertheless, many applications have normalized both shocks because it can be empirically difficult to identify both variances. We follow this practice. See Ursu et al. (2023) for more discussion.
Let vector $\boldsymbol{x}_{i}^{(cons)}$ collect the attributes of consumer $i$ (e.g., demographics, purchase history). These attributes affect the utility of an outside option as follows, with $\varepsilon_{i0}\sim N(0,1)$.
$u_{i0}=\eta_{0}+\boldsymbol{\eta}^{\prime}\boldsymbol{x}_{i}^{(cons)}+\varepsilon_{i0}.$
Each consumer has one free search. Afterwards, a search cost is incurred for each additional search. We allow some product-level attributes that affect search cost but not utility, collected in $\boldsymbol{x}_{ij}^{(ads)}$. We will refer to them as advertising or marketing attributes. One example of such attributes is the product rankings on the webpage. The search cost $c_{ij}$ is specified as
$\log(c_{ij})=\alpha_{0}+\boldsymbol{\alpha}^{\prime}\boldsymbol{x}_{ij}^{(ads)}.$
We use $d^{(prod)}$, $d^{(ads)}$, and $d^{(cons)}$ to denote the numbers of attributes in $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ , and $\boldsymbol{x}_{i}^{(cons)}$, respectively. Let $d\equiv d^{(prod)}+d^{(ads)}+d^{(cons)}$ denote the number of all attributes. Let $\boldsymbol{\theta}=(\boldsymbol{\beta}^{\prime},\boldsymbol{\eta}^{\prime},\boldsymbol{\alpha}^{\prime},\eta_{0},\alpha_{0})^{\prime}$ denote our search model’s parameter. The dimension of $\boldsymbol{\theta}$ is $d+2$.
The search model predicts: (i) which products the consumer will search, (ii) which products the consumer will buy, and (iii) the order by which the consumer searches the products. We shall not delve into the details of deriving these predictions. It suffices to mention that the predictions are usually calculated using the optimal decision rule in Weitzman (1979).
It is useful to note that the search model can also be applied to contexts outside online browsing and purchase. For example, increasingly available geo-location data allow us to study consumers’ search behaviors in offline shopping. Shop visits and subsequent purchases can be modeled by a sequential search model. An interesting aspect is that the distances from shops to elevators or entrances may affect search costs (and thus be considered as a marketing attribute).
2.2 The data
Next, we describe the typical data used to estimate the search model. In doing so, we also make clear the range of data to which our pretrained NNE is applicable.
Continuing with the notation so far, we use $i$ to index consumers and $j$ to index products. Let $\boldsymbol{x}_{ij}^{(prod)}$ collect product attributes, $\boldsymbol{x}_{ij}^{(ads)}$ collect advertising attributes, and $\boldsymbol{x}_{i}^{(cons)}$ collect consumer attributes. A dataset ${\cal D}$ can be written as
${\cal D}$ =$=\{\boldsymbol{x},\;\boldsymbol{y}\},$ (1)
[…]
$\quad\boldsymbol{y}=\{\{\boldsymbol{y}_{ij}\}_{j=1}^{J}\}_{i=1}^{n}.$
Here, vector $\boldsymbol{y}_{ij}$ collects two outcome variables:
  • $y_{ij}^{(search)}$ is a dummy indicating whether the option $j$ is searched;
  • $y_{ij}^{(buy)}$ is a dummy indicating whether $j$ is bought.
In some applications, the order of searches is also available, so $\boldsymbol{y}_{ij}$ can have a third element $y_{ij}^{(order)}$. However, because the search order is not always available, in this paper we choose not to make use of search order information (so $y_{ij}^{(order)}$ is not required). Conceptually, it is possible to pretrain a NNE that uses the search order information.
Our pretrained NNE assumes that $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ and $\boldsymbol{x}_{i}^{(cons)}$ are all standardized to have zero means and unit variances. The means and variances here are computed across $i$ and $j$, e.g., the mean for $\boldsymbol{x}_{ij}^{(prod)}$ is $\frac{1}{nJ}\sum_{i=1}^{n}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$. If the attributes are not standardized, we standardize them before using the pretrained NNE to estimate $\boldsymbol{\theta}$. We then rescale the estimate of $\boldsymbol{\theta}$ to make it interpretable with respect to the original, non-standardized attributes.
Table 1: Specified Ranges for Dataset Size
$d^{(prod)}$ $d^{(ads)}$ $d^{(cons)}$ $J$ $n$
[…]
The size of ${\cal D}$ is determined by the numbers of attributes ($d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$), number of products per consumer ($J$), and number of consumers ($n$). These numbers vary across applications. Our current version of pretrained NNE can be applied to datasets of sizes within the ranges specified in Table 1. These ranges are configurable before pretraining.
A note about Table 1 is that no cap is imposed on $n$ for applying the pretrained NNE. But a cap on $n$ is necessary in pretraining, because we cannot simulate data of infinite size. This cap is currently set at $42e$ 4. Pretrained NNE is not directly applicable to data larger than this cap. However, if an application has larger data, we can split the data and average the estimates across the splits. [^2] The splitting and averaging are incorporated in our code at pnnehome.github.io.
Too few searches or purchases in data indicate difficulty for estimating $\boldsymbol{\theta}$. Therefore, we do not pretrain NNE to accommodate such data. Specifically, we look at:
  • Buy rate: fraction of consumers that bought some product (not outside good).
  • Search rate: fraction of consumers that did at least one non-free search.
We require that the buy rate is at least 0.5% and the search rate is at least 1%. Again, these thresholds are configurable prior to pretraining. During pretraining, simulated datasets where search rates or buy rates are lower than thresholds are dropped.
As will be described later, the pretraining includes simulated datasets with non-normally distributed attributes. Nevertheless, the pretrained NNE may still perform poorly if certain attributes are too heavily skewed or have too many outliers. Therefore, we recommend using the usual practices to prepare data as in any empirical analysis. These include trimming, winsorizing, transforming variables, and dropping any attribute that causes (almost) collinearity.
3 Pretrained NNE
This section describes the construction of pretrained NNE. Although we aim to make the description self-contained, more emphasis is given to where the pretrained NNE differs from the standard NNE (Wei and Jiang 2024).
3.1 Overview
Let ${\cal D}$ denote a generic dataset as defined in expression (1). Let $\boldsymbol{\theta}$ collect the parameters of the search model specified in Section 2.1. The task of NNE is to map ${\cal D}$ to an estimate of $\boldsymbol{\theta}$. It does so through two functions, $\boldsymbol{f}$ and $\boldsymbol{h}$.
$\widehat{\boldsymbol{\theta}}=\boldsymbol{f}\big{[}\>\underbrace{\boldsymbol{h}({\cal D})}_{\boldsymbol{m}}\>\big{]}.$
  • The first function $\boldsymbol{h}$ calculates a set of data patterns of ${\cal D}$, collected in vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$. This function is specified by us (rather than trained by examples).
  • The second function $\boldsymbol{f}$ is a neural net that maps from the data patterns to an estimate of $\boldsymbol{\theta}$. We write $\widehat{\boldsymbol{\theta}}=\boldsymbol{f}(\boldsymbol{m})$.
Conceptually, one could forgo specifying $\boldsymbol{h}$ and instead train a neural net mapping directly from ${\cal D}$ to $\boldsymbol{\theta}$. But the size of ${\cal D}$ is much larger than $\boldsymbol{m}$. In addition, the mapping from ${\cal D}$ to $\boldsymbol{\theta}$ tends to be much more complex than from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. So the specification of function $\boldsymbol{h}$ substantially reduces the complexity of the learning task. The use of aggregate data patterns also allows privacy-preserving estimation, as discussed in Section 1.1.
In general, the vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ may include any data pattern of ${\cal D}$. The foremost candidates are various data moments (this is why we use notation $\boldsymbol{m}$), such as means, variances, correlations. More sophisticated candidates are coefficients of reduced-form regressions. The idea of using reduced-form coefficients as “moments” originated from indirect inference (Gourieroux et al. 1993). Including reduced-form coefficients in $\boldsymbol{m}$ turns out to substantially improve the estimation accuracy of NNE, and is an important extension to Wei and Jiang (2024).
The detailed specification of $\boldsymbol{h}$ in our pretrained NNE is given in appendix. Below, we give a summary.
  • First included in $\boldsymbol{m}$ are data size and summary statistics: $d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$, $n$, $J$, buy rate, search rate, the average number of searches per consumer.
  • To capture search decisions, we include the coefficients from logit regression that predicts $y_{ij}^{(search)}$. Similarly, to capture purchase decisions, we include the coefficients from multinomial logit regression that predicts $y_{ij}^{(buy)}$ conditional on search, i.e., $y_{ij}^{(search)}=1$. These regressions are at the product level.
  • We include the coefficients from a set of regressions similar to above, but at the consumer level (e.g., whether consumer $i$ bought anything, whether $i$ made any non-free search).
Next, we turn to the function $\boldsymbol{f}$, which is the core of the pretrained NNE. It is a neural net trained using simulated datasets. The training is conducted as follows.
  • We use the search model to simulate $L$ datasets ${\cal D}^{(1)}$, ${\cal D}^{(2)}$, …, ${\cal D}^{(L)}$, each under a different parameter, $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$.
  • For each $\ell=1,2,...,L$, we compute the data patterns $\boldsymbol{m}^{(\ell)}=\boldsymbol{h}({\cal D}^{(\ell)})$.
  • Each pair $\{\boldsymbol{m}^{(\ell)},\boldsymbol{\theta}^{(\ell)}\}$ is a training example. We use $\{\boldsymbol{m}^{(\ell)},\boldsymbol{\theta}^{(\ell)}\}_{\ell=1}^{L}$ to train a neural net to predict $\boldsymbol{\theta}$ from $\boldsymbol{m}$.
We have left out three details in this training process: (i) the configuration of the neural net, (ii) the distribution from which we draw $\boldsymbol{\theta}^{(\ell)}$, and (iii) the distribution from which we draw the product, consumer, and advertising attributes in ${\cal D}^{(\ell)}$. These details are discussed in Section 3.2. At a high level, the distributions in (ii) and (iii) should be broad so that NNE can accommodate a variety of real datasets in applications, and yet not too broad so that NNE wastes resources learning scenarios that are unlikely to arise in reality. A good balance makes the training more effective for any given $L$. Meanwhile, it is important for us to point out the pretrained NNE’s estimates in applications (Section 4) are not sensitive to the specifications of details (i)-(iii).
The description so far has focused on point estimates. For the standard error of $\widehat{\boldsymbol{\theta}}$, we recommend bootstrapping. This boils down to applying the pretrained NNE to resampled copies of ${\cal D}$. Because it incurs very little cost to apply pretrained NNE to each copy of data, bootstrapping standard errors is not computationally expensive at all. [^3] The bootstrapping here omits the uncertainty from the training of neural net. However, because of the large number of training examples ($L\simeq 1$ million), this uncertainty is small relative to the sampling uncertainty of ${\cal D}$.
A final issue in real applications is that some datasets may not be suitable for our pretrained NNE, because their true data generating processes are poorly described by the search model in Section 2.1. Actually, in such cases it is advisable not to apply this search model. For this reason, it will be useful to have a way to detect ill-suited datasets. We describe a detection method in the appendix. The basic idea is to check whether the real data patterns are an outlier compared to the data patterns used in pretraining.
3.2 Pretraining details
So far we have given an overview of how the pretrained NNE is constructed. Next, we provide more details on the training of the neural net $\boldsymbol{f}$. Readers more interested in the applications may skip these details without much loss of continuity.
3.2.1 The neural net
A neural net accepts fixed-length input and gives fixed-length output. But the search model parameter $\boldsymbol{\theta}$ has a dimension of $d+2$, where $d$ is the total number of attributes and it varies across different datasets. We address this issue through padding. Specifically, recall that $d\equiv d^{(prod)}+d^{(cons)}+d^{(ads)}$. Let subscript “max” denote the maximal value of a dimension. For example, Table 1 sets $d_{max}^{(prod)}=8$ and $d_{max}=8+5+2=15$. We configure the neural net to always output $d_{max}+2$ values. Out of these output values, only $d+2$ values are useful. The rest are the padded “idle” values. Using ellipses to represent idle values, we can write the output as:
$(\enspace\alpha_{0}\enspace,\enspace\boldsymbol{\alpha}^{\prime}\enspace,\underbrace{\enspace\cdot\cdot\cdot\enspace}_{d_{max}^{(ads)}-d^{(ads)}},\enspace\eta_{0}\enspace,\enspace\boldsymbol{\eta}^{\prime}\enspace,\underbrace{\enspace\cdot\cdot\cdot\enspace}_{d_{max}^{(cons)}-d^{(cons)}},\enspace\boldsymbol{\beta}^{\prime}\enspace,\underbrace{\enspace\cdot\cdot\cdot\enspace}_{d_{max}^{(prod)}-d^{(prod)}}).$
In training, the loss function ignores idle output values. In applications, the idle output values are disregarded.
A similar arrangement applies to the neural net’s input. The length of $\boldsymbol{m}$ varies by the numbers of attributes in the data. For our current settings in Table 1, the maximal length of $\boldsymbol{m}$ is 164. Any shorter $\boldsymbol{m}$ is padded with zeros to this maximal length before we feed it into the neural net. The exact detail is given in the specification of $\boldsymbol{m}$ in appendix.
We currently use $L\simeq$ 1e6 training examples (this is after dropping “corner” datasets with very low search or buy rates; see Section 2.2). We hold out $45e$ 4 examples for validation. Increasing $L$ further only marginally reduces the validation loss. We also use hold-out validation to choose hyperparameters of the neural net. One particular hyperparameter is the neural net architecture. We end up using three hidden layers with 512, 256, and 256 nodes. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the neural net architecture.
3.2.2 The parameter prior
When generating the training datasets, we draw $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$ from a distribution of $\boldsymbol{\theta}$. As will be explained in Section 3.3, this distribution of $\boldsymbol{\theta}$ effectively plays the role of a prior for $\boldsymbol{\theta}$. We want the prior to be sufficiently broad to cover possible values of $\boldsymbol{\theta}$ in real applications, but not too broad so that many training datasets become unrealistic and thus wasteful. Below, we explain the prior used in our pretrained NNE. We have also tried several alternative priors. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the prior specification.
Unlike Bayesian estimation on a single dataset, a unique challenge in specifying our prior is that we need to accommodate datasets with varying dimensions ($d^{(prod)}$, $d^{(cons)}$, $d^{(ads)}$). Consider key statistics such as buy rate and search rate. In real applications, there is no inherent reason for us to expect these key statistics to vary systematically across datasets of different dimensions. Therefore, among training datasets, we would like the key statistics not to vary systematically by data dimensions, either.
This desire rules out common priors such as $\beta_{k}\sim N(0,1)$ for $k=1,...,d^{(prod)}$. With this prior, the buy rates would be systematically higher for simulated datasets with larger $d^{(prod)}$. To see why, note this prior implies $\lVert\boldsymbol{\beta}\rVert^{2}\sim\chi_{d^{(prod)}}^{2}$ which has an expectation equal to $d^{(prod)}$. Consider a simple case where $\boldsymbol{x}_{ij}^{(prod)}\sim N(\boldsymbol{0},\boldsymbol{I})$. Then, the variance of $\boldsymbol{\beta}^{\prime}\boldsymbol{x}_{ij}^{(prod)}$ for any given $\boldsymbol{\beta}$ equals $\lVert\boldsymbol{\beta}\rVert^{2}$, which is expected to increase with $d^{(prod)}$. A larger variance makes it more likely for at least one product to be better than outside option, leading to a higher buy rate.
To this end, we specify a prior where $\lVert\boldsymbol{\beta}\rVert$ is invariant to $d^{(prod)}$:
$\boldsymbol{\beta}\sim\sigma\cdot\chi_{1}\cdot\boldsymbol{\nu},$ (2)
where $\sigma$ is a constant, $\chi_{1}$ is chi distributed, and $\boldsymbol{\nu}$ is a unit vector with a uniform distribution on the unit sphere. An implication of this prior is $\lVert\boldsymbol{\beta}\rVert^{2}\sim\sigma^{2}\cdot\chi_{1}^{2}$, which is invariant to $d^{(prod)}$. We use the same type of prior for $\boldsymbol{\eta}$ and $\boldsymbol{\alpha}$.
The prior (2) can also be understood as letting the total impact of product attributes (measured by $\lVert\boldsymbol{\beta}\rVert$; recall that all attributes are standardized) be invariant to $d^{(prod)}$. This means that when there are more product attributes, we expect that each individual attribute has a lower chance to have a large effect on utility. This expectation seems to align with what we see in real applications, where typically only a few “key” product attributes have substantial effects on consumers’ search and purchase decisions.
The discussion so far has left out the priors for the intercepts $\alpha_{0}$ and $\eta_{0}$, which directly affect search cost and outside utility. We specify the priors of these intercepts as normally distributed. Their means and variances are specified to generate realistic search rates and buy rates in training datasets (neither too small nor too close to 1).
3.2.3 Attributes ($\boldsymbol{x}$)
To generate a training dataset ${\cal D}^{(\ell)}$, we first draw $\boldsymbol{x}^{(\ell)}$ from a distribution of $\boldsymbol{x}$. Next, given $\boldsymbol{x}^{(\ell)}$ and $\boldsymbol{\theta}^{(\ell)}$, we use the search model to simulate $\boldsymbol{y}^{(\ell)}$. Finally, we let ${\cal D}^{(\ell)}=\{\boldsymbol{x}^{(\ell)},\boldsymbol{y}^{(\ell)}\}$. This procedure requires us to specify a distribution of $\boldsymbol{x}$. Conceptually, this distribution should be broad enough to cover common values of $\boldsymbol{x}$ in real applications, but not too broad so that many training datasets become unrealistic and thus wasteful. Below, we describe the distribution of $\boldsymbol{x}$ currently used by us. It is important to note that the pretrained NNE’s estimates in applications (Section 4) are not sensitive to how we specify this distribution.
First, we draw the size of $\boldsymbol{x}$. The size is determined by $d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$, $J$, and $n$. All these numbers are picked randomly from their respective ranges to which we would like the pretrained NNE to be applicable. Our current choices of these ranges are given in Table 1 (in Section 2.2). A special note is for $n$. We cap $n$ in pretraining (currently at 2e4), because we cannot simulate a dataset of infinite size. However, this cap does not restrict the applications of pretrained NNE, because we can always split a larger dataset and average the estimates across the splits.
Next, we draw the attributes in $\boldsymbol{x}$. There are three features of $\boldsymbol{x}$ commonly seen in real applications that we would like to incorporate. The first feature is that many attributes show substantial consumer-level variations. For example, some consumers see prices systematically higher than other consumers. Therefore, we draw $\boldsymbol{x}$ with two components: a within-consumer component plus a cross-consumer component (the latter is also known as consumer fixed effects). An exception here is $\boldsymbol{x}_{i}^{(cons)}$, which by definition only has the cross-consumer component.
The second feature is that individual attributes often follow non-normal distributions. To this end, we draw each attribute initially as normal and then transform it into one of the following non-normal distributions: (i) a dummy, (ii) a scale from 1 to 5, and (iii) a bell-shaped but skewed continuous distribution. In principle, more types of distributions could be included. However, when we do so, the pretrained NNE’s estimates in applications (Section 4) remain mostly unchanged. So these three types of distributions seem sufficient.
The third feature is that attributes are often correlated. To this end, we let the initial attribute draws follow multivariate normal distributions with randomly drawn correlation matrices. The correlations persist through subsequent transformation into non-normal distributions. For example, if two attributes are positively correlated when first drawn as normal, they will remain positively correlated when one of them is transformed into a dummy.
3.3 Link to Bayesian estimation
The eventual test for a pretrained model is to see how it performs in applications. We do so for the pretrained NNE in Section 4 and find that it provides accurate and sensible estimates. That said, from a theoretical standpoint, it is useful to understand the econometric meaning of the estimate by pretrained NNE. Below, we will explain that pretrained NNE has a familiar Bayesian interpretation, in that it calculates the posterior mean of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. In words, it provides the best estimate of the structural model’s parameter $\boldsymbol{\theta}$ given the data patterns $\boldsymbol{m}$.
Continuing with the notation in (1), let ${\cal D}=\{\boldsymbol{x},\boldsymbol{y}\}$ denote a dataset. As before, $\boldsymbol{x}$ and $\boldsymbol{y}$ here collect the observations across all consumers. The vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ denotes a set of data patterns. We start with a result with the standard (non-pretrained) NNE, studied in Wei and Jiang (2024). The main difference between standard NNE and pretrained NNE is that the former is trained for a specific real dataset. The training examples do not simulate $\boldsymbol{x}$ but directly use the observed $\boldsymbol{x}$ from that real dataset.
The standard NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ as $L\rightarrow\infty$ (see Wei and Jiang 2024). Note that the result does not require large $n$ and holds for any value of $n$. Also note that we have made it explicit that the expectation conditions on a specific $\boldsymbol{x}$ [^4] The notations of Wei and Jiang (2024) omit the conditioning on $\boldsymbol{x}$ because the paper assumes an underlying probability space that is already conditional on the observed $\boldsymbol{x}$. The $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ is known as a limited-information Bayesian posterior mean (e.g., Kim 2002), where “limited-information” indicates that the posterior relies on less information than the full data. To compare, the full-information posterior is $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{y},\boldsymbol{x})$. Ideally, we would like to compute the full-information posterior, but when doing so is difficult, the limited-information posterior provides a practical alternative.
We now turn to the pretrained NNE. In this case, the training examples do not assume a specific $\boldsymbol{x}$, but instead simulate $\boldsymbol{x}$ from a distribution (as described in Section 3.2). Consequently, the estimate by the pretrained NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ as $L\rightarrow\infty$. Importantly, note that the expectation no longer conditions on $\boldsymbol{x}$.
Given this convergence result, a natural question is how to interpret $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. Formally, it is another limited-information posterior mean. For a more intuitive interpretation, consider a practical scenario where a researcher wants to apply the search model to a company’s data. For privacy reasons, the company provides the researcher with only aggregate data patterns $\boldsymbol{m}$, instead of the full data $\{\boldsymbol{x},\boldsymbol{y}\}$. In this case, our best estimate for the search model parameter will be precisely $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. By definition,
[…]
In equation (3), the term $P(\boldsymbol{m}|\boldsymbol{\theta},\boldsymbol{x})$ is defined by the search model, i.e., the likelihood of observing the data patterns $\boldsymbol{m}$ given $\boldsymbol{\theta}$ and $\boldsymbol{x}$. The term $P(\boldsymbol{\theta},\boldsymbol{x})$ must be specified by the researcher as a prior. Typically, the prior for $\boldsymbol{\theta}$ is specified independently of $\boldsymbol{x}$: $P(\boldsymbol{\theta},\boldsymbol{x})=P(\boldsymbol{\theta})P(\boldsymbol{x})$. These two components of the prior, $P(\boldsymbol{\theta})$ and $P(\boldsymbol{x})$, correspond to the distributions we use to draw $\boldsymbol{\theta}$ and $\boldsymbol{x}$ during the pretraining (described in Section 3.2).
This link to Bayesian estimation guides our specification of $\boldsymbol{m}$. Because the pretrained NNE computes $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$, $\boldsymbol{m}$ should include the data patterns that are informative about $\boldsymbol{\theta}$. Meanwhile, we would like the computation of $\boldsymbol{m}$ to be inexpensive. These considerations have led us to use the coefficients from various reduced-form regressions of $\boldsymbol{y}$ on $\boldsymbol{x}$ (see Section 3.1).
4 Applications
We evaluate the pretrained NNE’s performance with 12 real datasets. In particular, we compare it to SMLE, which is the prevailing approach to estimate search models. Below, we first briefly describe each dataset. Then, we conduct Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices. Finally, we estimate the search model on the real data entirely.
Before proceeding, we briefly describe existing approaches to estimate search models, for readers unfamiliar with the literature. Standard SMLE with accept-reject simulations is infeasible for search models because it requires prohibitive computational costs. A common workaround in literature is to smooth the simulated likelihood function (e.g., Honka and Chintagunta 2016, Ursu 2018). However, estimates are sensitive to smoothing factors and tuning these factors requires costly Monte Carlo experiments. More recently, Jiang et al. (2021) and Chung et al. (2023) extend the Geweke-Hajivassiliou-Keane (GHK) simulation algorithm to the likelihood of search models. The GHK-based SMLE does not require smoothing, but is much more complex to implement. We will consider both the smoothed SMLE and the GHK-based SMLE.
4.1 Datasets
Our 12 datasets come from four different settings (an anonymous online retailer, Expedia, JD, and Trivago). Table 2 provides summary statistics. We see that these datasets vary substantially in the number of search sessions, numbers of attributes, buy rate, and search rate. So they cover a fairly wide range of real-world scenarios.
For ease of presentation, the estimates reported in this section assume that each attribute is mean-zero with a standard deviation of 0.2 (instead of 1). The 0.2 standard deviation makes coefficients ($\boldsymbol{\beta}$, $\boldsymbol{\eta}$, $\boldsymbol{\alpha}$) larger and thus more visible in figures. For estimation, all attributes are standardized to unit variance before being fed into the pretrained NNE (see Section 2.2).
We briefly describe each dataset below.
Table 2: Summary Statistics of 12 Datasets
$n$ $d^{(prod)}$ $d^{(cons)}$ $d^{(ads)}$ buy rate search rate
eCommerce: laptop 10,000 6 0 0 5.9% 21.5%
eCommerce: vacuum 10,000 4 0 0 8.5% 19.0%
eCommerce: washer 10,000 5 0 0 7.8% 22.1%
Expedia: destination 1 1,258 6 2 1 11.5% 8.7%
Expedia: destination 2 897 6 2 1 3.8% 6.0%
Expedia: Cancun 2,135 6 0 1 1.3% 30.8%
Expedia: Manhattan 1,369 6 0 1 3.1% 28.7%
JD: desktop channel 5,896 4 2 0 16.4% 26.5%
JD: mobile channel 4,069 4 2 0 11.2% 12.2%
JD: app channel 10,000 4 2 0 8.0% 12.9%
Trivago: desktop channel 2,082 7 0 1 2.8% 40.0%
Trivago: mobile channel 2,726 7 0 1 2.5% 29.1%
Notes: $n$ is the number of consumers or search sessions. $d^{(prod)}$, $d^{(cons)}$, and $d^{(ads)}$ are the numbers of product attributes, consumer attributes, and advertising attributes, respectively. Buy rate is the percentage of consumers who made a purchase. Search rate is the percentage of consumers who made at least one non-free search.
(1) eCommerce: laptop
Our first dataset consists of consumer searches and purchases in an anonymous large online store, provided by the marketing company REES46. The data is available on Kaggle.com. [^5] https://www.kaggle.com/datasets/mkechinov/ecommerce-behavior-data-from-multi-category-store The store sells products across many categories. This dataset focuses on laptops. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and five brand dummies: Asus, HP, Lenovo, MSI, and Xiaomi (the baseline brand is Acer). There are no consumer or advertising attributes. There are a very large number of search sessions. To ease the computational burden of SMLE, we take a subsample of 10,000 search sessions.
(2) eCommerce: vacuum
This dataset includes the vacuum category in the anonymous online store. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and three brand dummies: Philips, Samsung, and Vitek (the baseline brand is Bosch). We again take a subsample of 10,000 search sessions.
(3) eCommerce: washer
This dataset includes the washer category in the anonymous online store. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and four brand dummies: Beko, Candy, Indesit, Midea (the baseline brand is Atlant). We again take a subsample of 10,000 search sessions.
(4) Expedia: destination 1
The International Conference on Data Mining (ICDM 2013) organized a contest to predict consumer searches and bookings for hotels on Expedia. The data for the contest is publicly available on Kaggle.com. [^6] https://www.kaggle.com/competitions/expedia-personalized-sort/overview It has been used in several papers to study consumer search behaviors (Ursu 2018, Greminger 2022, Compiani et al. 2023). The data includes multiple travel destinations. “Destination 1” refers to the largest destination. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include star rating, review score, a chain indicator, location score, price, and promotion. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The advertising attribute $\boldsymbol{x}_{ij}^{(ads)}$ includes the ranking of the hotel.
(5) Expedia: destination 2
This dataset consists of the search sessions for the second largest destination in the Expedia data from ICDM 2013. There are only $n=897$ search sessions, slightly below the minimum threshold of $n$ set for the pretrained NNE (see Table 1). Nonetheless, we decide to include the dataset, as it allows us to observe how the pretrained NNE performs slightly outside its intended range of applications.
(6) Expedia: Cancun
The Analytics at Wharton Initiative also provides data on hotel searches and bookings on Expedia. [^7] https://analytics.wharton.upenn.edu/open-research-opportunities/predicting-optimizing-consumer-response-to-product-search-results/ The data is not openly available but can be accessed after a request with the Initiative. The data includes four travel destinations (Cancun, Manhattan, Paris and Budapest). Cancun is the largest destination.
(7) Expedia: Manhattan
This dataset consists of the search sessions for the second largest destination (Manhattan) in the Expedia data from Analytics at Wharton.
(8) JD: desktop channel
The 2020 MSOM Data Driven Research Challenge provided data on consumer experiences on JD.com. The data was made available to INFORMS members in 2019.888https://connect.informs.org/msom/events/datadriven2020 JD.com is China’s second largest business-to-consumer online retail platform. The data covers many aspects of consumer experiences on the platform, from website browsing to order fulfillment (see Shen et al. 2020 for details). We focus on consumer search and purchase activities. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price, a dummy for first-party sellers, and two unnamed attributes. There are no advertising attributes.
The JD.com data includes three channels: desktop, mobile, and app. This dataset focuses on the desktop channel.
(9) JD: mobile channel
This dataset includes the mobile channel from the JD.com data.
(10) JD: app channel
This dataset includes the app channel from the JD.com data. There are a large number of search sessions. To ease the computational burden of SMLE, we use a subsample of 10,000 search sessions.
(11) Trivago: desktop channel
The ACM RecSys Challenge 2019 provides consumer data on Trivago.com. [^9] https://recsys.acm.org/recsys19/challenge/ The data is not standard for search models because purchases are not very well defined in this context, as explained below. Nonetheless, we decide to include the data, because it allows us to observe how the pretrained NNE performs somewhat outside the intended context of the search model.
More specifically, Trivago is a hotel meta-search engine that aggregates information from hotels and travel websites (similar to how Google Flights aggregates flight information). On Trivago.com, consumers receive a list of hotels. They can click on any hotel to see more information. They can also click a “View Deal” button that directs them to the external site for possible booking. But what happens at external sites is not observed in the data. We treat a click for more information as a search and a “View Deal” click as a purchase. Sometimes a consumer clicks “View Deal” for multiple hotels, in which case we randomly pick one as the purchase. Cases where consumers click “View Deal” without first clicking for more information are dropped (about 5%).
The Trivago data includes search sessions from both the desktop channel and mobile channel. This dataset focuses on the desktop channel.
(12) Trivago: mobile channel
This dataset consists of the search sessions from the mobile channel in the Trivago data.
4.2 Monte Carlo studies
Table 3: Estimates in Monte Carlo Study with eCommerce: Laptop
Truth Pretrained NNE SMLE
SMLE $R=50$
SMLE
SMLE
$R=50$ $\boldsymbol{\lambda}=(4,2)$
$\alpha_{0}$5.0 -5.009 (0.006) -5.779 (0.010) -4.696 (0.023) -3.347 (0.011) -2.767 (0.011)
$\eta_{0}$ 4.8 4.797 (0.002) 4.732 (0.004) 4.851 (0.009) 4.981 (0.005) 3.999 (0.006)
$\beta_{1}$0.6 -0.603 (0.002) -0.579 (0.003) -0.375 (0.005) -0.457 (0.005) -0.366 (0.005)
$\beta_{2}$0.5 -0.488 (0.002) -0.485 (0.003) -0.300 (0.005) -0.370 (0.005) -0.295 (0.005)
$\beta_{3}$0.3 -0.292 (0.002) -0.290 (0.003) -0.188 (0.005) -0.228 (0.006) -0.183 (0.005)
$\beta_{4}$1.0 -1.010 (0.003) -0.968 (0.003) -0.627 (0.005) -0.771 (0.005) -0.616 (0.005)
$\beta_{5}$0.4 -0.399 (0.003) -0.390 (0.003) -0.255 (0.005) -0.315 (0.006) -0.250 (0.005)
$\beta_{6}$0.5 -0.489 (0.003) -0.479 (0.003) -0.316 (0.006) -0.392 (0.006) -0.312 (0.005)
Notes: $R$ is the number of simulations per consumer in SMLE. $\boldsymbol{\lambda}$ is the smoothing factors in the smoothed SMLE. Results are averaged across 100 Monte Carlo repetitions. Numbers in parentheses are the standard errors for the averages.
Our first set of tests are Monte Carlo studies. Specifically, for each of the 12 real datasets, we: (i) keep the product, consumer, and advertising attributes, but (ii) replace the search and purchase choices with those simulated by the search model under a “true” $\boldsymbol{\theta}$. Then, we try to recover $\boldsymbol{\theta}$ from this dataset. Knowing the ground truth of $\boldsymbol{\theta}$ allows us to measure estimation accuracy. The true $\boldsymbol{\theta}$ is set to be close to its estimate later in Section 4.3.
Figure 2: Monte Carlo Study with eCommerce: Laptop
Notes: Results are based on 100 Monte Carlo repetitions.
We first closely examine the results from the eCommerce: laptop dataset, reported in Table 4.2. Column 1 shows the true value of $\boldsymbol{\theta}$. Column 2 shows the estimate by the pretrained NNE. Columns 3 - 6 show the estimates by SMLE. The $R$ denotes the number of simulations per consumer in SMLE and $R=50$ is typical in literature. For the smoothed SMLE, $\boldsymbol{\lambda}$ collects the smoothing factors for search choices and for purchase choices. [^10] Some applications use the same smoothing factor for search and purchase choices. But doing so here worsens RMSE substantially. Whether multiple smoothing factors are necessary is generally application-dependent. The table shows three settings of $\boldsymbol{\lambda}$ to illustrate the sensitivity of SMLE’s estimate to $\boldsymbol{\lambda}$. We actually have conducted an expensive grid search of $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}=(4,1)$ gives the smallest root mean square error (RMSE) in this application.
The upper plot of Figure 2 reports the RMSEs of the different estimators. The pretrained NNE has the lowest RMSE, outperforming even the SMLE that uses GHK and a very large $R=1000$. Although we generally expect MLE to be asymptotically efficient, the performance of SMLE is affected by simulation errors. In search models, simulating the likelihood is particularly difficult because of the large number of possible search and purchase combinations. Our result here suggests that this likelihood simulation remains challenging despite workarounds such as smoothing and GHK. The result also shows that pretrained NNE can achieve high estimation accuracy despite being off-the-shelf.
The lower plot of Figure 2 reports the computation time needed for a single estimation run. All the times are measured on the same hardware. Recall that this dataset is a subsample from a single product category of the original multi-category data (see Section 4.1). [^11] In the original data, there are 624 categories. The laptop category has 41,918 search sessions. Our dataset is a subsample of 10,000 search sessions. Even so, the SMLE requires between 1.5 and 10 hours. The pretrained NNE takes about 1 second (so the orange bar is barely visible). This represents a time saving of 3 to 4 orders of magnitude. In addition, note that implementing SMLE requires additional costs not reflected in Figure 2, such as tuning smoothing factors, tuning optimization routine, trying starting points. Each of these tasks requires multiple estimation runs. Moreover, Figure 2 does not reflect the coding cost of SMLE, which can be substantial with GHK.
Next, we expand the examination to all 12 datasets. The results are reported in Table 4. The table omits the smoothed SMLE because, as we have seen, its estimates are sensitive to the smoothing factors $\boldsymbol{\lambda}$ (and the optimal $\boldsymbol{\lambda}$ is application-dependent). The main finding with Table 4 is qualitatively the same as what we have seen with the eCommerce: laptop dataset. The pretrained NNE achieves either comparable or better estimation accuracy at drastically lower computational costs.
It is worth noting that the sizes of the 12 datasets here are not large (see Table 2). In fact, recall that each dataset focuses on only a subset (a single product category, travel destination, or sales channel) of its source. Further, four of the 12 datasets have been subsampled to reduce the computation burden. Because computation time scales with data size, the absolute time savings on large datasets will be much greater than those indicated by Table 4.
Table 4: Monte Carlo Studies Across 12 Datasets
Pretrained NNE SMLE (GHK) $R=50$ SMLE (GHK) $R=1000$
RMSE
eCommerce: laptop 0.094 (0.003) 0.795 (0.010) 0.453 (0.009)
eCommerce: vacuum 0.095 (0.003) 0.352 (0.005) 0.176 (0.005)
eCommerce: washer 0.179 (0.009) 0.614 (0.007) 0.361 (0.008)
Expedia: destination 1 0.451 (0.012) 0.591 (0.022) 0.501 (0.015)
Expedia: destination 2 0.748 (0.023) 0.964 (0.035) 0.833 (0.028)
Expedia: Cancun 0.485 (0.021) 1.814 (0.055) 1.083 (0.045)
Expedia: Manhattan 0.500 (0.018) 1.449 (0.054) 0.855 (0.040)
JD: desktop channel 0.175 (0.005) 0.392 (0.008) 0.256 (0.008)
JD: mobile channel 0.234 (0.007) 0.375 (0.011) 0.273 (0.009)
JD: app channel 0.161 (0.006) 0.335 (0.006) 0.200 (0.008)
Trivago: desktop channel 0.423 (0.016) 2.462 (0.050) 1.521 (0.044)
Trivago: mobile channel 0.422 (0.021) 1.067 (0.037) 0.660 (0.030)
Compute time (mins)
eCommerce: laptop 0.02 132.72 630.06
eCommerce: vacuum 0.02 59.42 371.52
eCommerce: washer 0.02 126.48 552.85
Expedia: destination 1 0.01 20.35 126.20
Expedia: destination 2 0.01 17.53 94.03
Expedia: Cancun 0.02 27.13 151.97
Expedia: Manhattan 0.02 19.01 101.17
JD: desktop channel 0.02 50.43 306.65
JD: mobile channel 0.01 33.53 213.45
JD: app channel 0.02 85.27 495.20
Trivago: desktop channel 0.02 33.98 212.17
Trivago: mobile channel 0.01 46.50 281.87
Notes: $R$ is the number of simulations per consumer in SMLE. $R=50$ is typical in literature. Results are averaged across 100 Monte Carlo repetitions. Numbers in parentheses are the standard errors for the averages.
4.3 Real-data estimates
Figure 3: Real-Data Estimates Across 12 Datasets
eCommerce: laptop
eCommerce: vacuum
eCommerce: washer
Expedia: destination 1
Expedia: destination 2
Expedia: Cancun
Expedia: Manhattan
JD: desktop
JD: mobile
JD: app
Trivago: desktop
Trivago: mobile
Our second set of tests estimates the search model entirely from real data – that is, both real attributes and real consumer choices. We no longer have a ground truth for $\boldsymbol{\theta}$ and cannot directly measure estimation accuracy. Therefore, we assess whether the pretrained NNE’s estimates are sensible, by comparing them with the estimates from the GHK-based SMLE that uses the very large $R=1000$.
Figure 3 shows the estimates across the 12 datasets. Each plot corresponds to one dataset. The horizontal axis lists the individual parameters in $\boldsymbol{\theta}$. The blue bars are estimates by SMLE and the orange bars are estimates by pretrained NNE. A 95% confidence interval is shown on each bar. Overall, the estimates by pretrained NNE and by SMLE are closely matched.
We omit reporting computation times in Figure 3, because they are quantitatively similar to those reported in Table 4 for Monte Carlo studies. Across all the datasets, the pretrained NNE uses only a fraction of the computation time required by SMLE.
5 Conclusion
This paper explores the feasibility of pretraining an off-the-shelf machine learning model to estimate a given structural econometric model. Importantly, we do not use the machine learning model to capture relationships in data. Instead, these relationships are prescribed by the structural model, and we use the machine learning model to recover this structural model’s parameter. Pretrained NNE enables users to estimate a structural model as easily as reduced-form regressions, while providing equal or better accuracy than existing estimation methods (at least in the context of the search model).
Structural estimation is very valuable in empirical research, e.g., it offers insights into micro-founded consumer preferences and enables counterfactual analyses. However, potentially wider applications have been hindered by the high costs in both researcher effort and computation time. Pretrained estimators help remove these barriers.
The ease of computation offered by pretrained NNE may facilitate new applications. One notable area is to integrate structural models with real-time algorithms, such as dynamic pricing, ranking algorithms, recommender systems, bandit experiments (e.g., Agrawal et al. 2019, Jain et al. 2024). Because structural models provide estimates of the latent underlying consumer preferences, these estimates can be fed into real-time algorithms to improve their decisions, e.g., what price to set, what items to recommend. Counterfactual analyses can be used to evaluate and provide feedback to the algorithms. However, structural estimation is time-consuming, which has made such applications difficult. Pretrained NNE offers a way to estimate structural models in real time, allowing more seamless integration of structural models into these algorithms.
The intention of this paper is to explore the feasibility of pretrained estimators, rather than develop a comprehensive estimator for search models. This leaves areas where the pretrained NNE can be further improved. We note two potential areas. The first area is to increase the richness of the structural model. The search model considered in this paper is a standard yet simple one. Future work may allow for attributes that simultaneously affect utility and search costs, or allow for richer unobserved heterogeneity in utility and search costs. These extensions will require us to find additional reduced-form patterns to help identify the more complex structural model.
A second area for improvement is to address potential endogeneity. For example, if researchers have an instrumental variable (IV) for price, how do we build a pretrained estimator that can leverage the instrument? At a high level, addressing this problem involves two tasks. The first task is to generate training datasets that include instruments. The second task is to pass relevant information regarding the instrument to the neural net. These tasks are not straightforward. However, incorporating IV in NNE will be a significant and valuable extension, because we frequently need to address endogeneity in structural estimation applications.
Appendix
Appendix A Data patterns
In Section 3.1 we have briefly described the data patterns $\boldsymbol{m}$ that feed into the neural net. Below, we provide the exact specification of $\boldsymbol{m}$.
We first recall some notation that will be needed to describe $\boldsymbol{m}$. Vectors $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$, and $\boldsymbol{x}_{i}^{(cons)}$ collect the product, advertising, and consumer attributes. It is assumed that all attributes are standardized. As before, let $d^{(prod)}$, $d^{(ads)}$, and $d^{(cons)}$ denote the numbers of attributes. These numbers are dataset-dependent. Let $d_{max}^{(prod)}$, $d_{max}^{(ads)}$, and $d_{max}^{(cons)}$ denote the maximum values that we currently set for the numbers of attributes (given in Table 1). These maximum values are not dataset-dependent.
We now define several consumer-level variables that will be used in $\boldsymbol{m}$. Recall that $J$ is the number of products per consumer and $n$ is the number of consumers.
  • Let $\bar{\boldsymbol{x}}_{i}^{(prod)}=\frac{1}{J}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$ denote the average product attributes for consumer $i$.
  • Let $\bar{\boldsymbol{x}}_{i}^{(ads)}=\frac{1}{J}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(ads)}$ denote the average advertising attributes for consumer $i$.
  • Let $\widetilde{y}_{i}^{(search)}=\sum_{j=1}^{J}y_{ij}^{(search)}$ denote the number of searches made by consumer $i$.
  • Let $\widetilde{y}_{i}^{(nf)}=\mathbb{I}\big{\{}\widetilde{y}_{i}^{(search)}>1\big{\}}$ indicates whether consumer $i$ did any non-free search.
  • Let $\widetilde{y}_{i}^{(buy)}=\sum_{j=1}^{J}y_{ij}^{(buy)}$, which is a dummy for whether consumer $i$ made a purchase.
The elements of $\boldsymbol{m}$ for any given dataset ${\cal D}$ are:
  1. 1.
    The buy rate, search rate, and average number of searches per consumer.
  2. 2.
    The values of $J$ and $n$.
  3. 3.
    A dummy vector indicating the numbers of product, advertising, and consumer attributes, shown as follows. Note that the vector has a fixed length of $d_{max}\equiv d_{max}^{(prod)}+d_{max}^{(ads)}+d_{max}^{(cons)}$.
    $(\quad\underbrace{1,...,1,}_{d^{(prod)}}\quad\underbrace{0,...,0,}_{d_{max}^{(prod)}-d^{(prod)}}\quad\underbrace{1,...,1,}_{d^{(ads)}}\quad\underbrace{0,...,0,}_{d_{max}^{(ads)}-d^{(ads)}}\quad\underbrace{1,...,1,}_{d^{(cons)}}\quad\underbrace{0,...,0}_{d_{max}^{(cons)}-d^{(cons)}}).$ (4)
  4. 4.
    The coefficients for $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $y_{ij}^{(search)}$ on $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\allowbreak\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(prod)},\bar{\boldsymbol{x}}_{i}^{(ads)}\}$. The non-intercept coefficients are padded with zeros and collected in a vector of fixed length $d_{max}$. The zero paddings enter the vector in the same way as in expression (4). This style of collecting regression coefficients applies to all the regressions below.
  5. 5.
    The coefficients in a multinomial logit regression of $y_{ij}^{(buy)}$ on $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)}\}$, conditional on $y_{ij}^{(search)}=1$. Intuitively, this regression describes how each consumer chooses among searched products.
  6. 6.
    The coefficients for $\{1,\;\boldsymbol{x}_{i}^{(cons)}\}$ in a linear regression of $\log(\widetilde{y}_{i}^{(search)})$ on $\{1,\;\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(ads)},\bar{\boldsymbol{x}}_{i}^{(prod)}\}$.
  7. 7.
    The coefficients for $\{1,\;\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $\widetilde{y}_{i}^{(nf)}$ on $\{1,\;\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(ads)},\bar{\boldsymbol{x}}_{i}^{(prod)}\}$.
  8. 8.
    The coefficients for $\{1,\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $\widetilde{y}_{i}^{(buy)}$ on $\{1,\;\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(ads)},\bar{\boldsymbol{x}}_{i}^{(prod)}\}$.
  9. 9.
    The means and standard deviations of the variables in all the regressions above, except for those with fixed values (e.g., the standard deviations of $\boldsymbol{x}_{ij}^{(prod)}$ always equal to 1).
All the coefficients above are estimated using ridge regression that imposes a small penalty on the coefficients. We find that using the ridge regression noticeably decreases the validation loss of the neural net (in predicting $\boldsymbol{\theta}$). We also find it helpful to use two separate penalty levels, and include the coefficients from both penalty levels in $\boldsymbol{m}$. Using two penalty levels further decreases the validation loss of the neural net. The two penalty levels we currently use are 0.001 and 1e-6. Using three or more penalty levels yields only marginal further decreases in the validation loss.
Appendix B Detecting ill-suited applications
For an off-the-shelf tool that estimates a structural model, it is helpful for users to have a way to detect whether a dataset is ill-suited for the structural model. A real dataset is ill-suited if the underlying data generating process is poorly described described by the assumed structural model. If we apply the pretrained NNE to an ill-suited dataset, the real data patterns will likely be outside the range of data patterns used in pretraining. As a result, the neural net will be extrapolating and the estimates may not be meaningful.
Thus, we want to detect whether the $\boldsymbol{m}$ based on a given real dataset is atypical compared to the training examples $\{\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$. This task is not straightforward because $\boldsymbol{m}$ can have hundreds of elements, and these elements are not independent. Even if each individual element of $\boldsymbol{m}$ appears typical on its own, the vector $\boldsymbol{m}$ can still be atypical if it contradicts a relationship between multiple elements that holds in the training examples.
We propose a practical way to detect atypical $\boldsymbol{m}$, by making use of a tree-based model (boosted tree or random forest). In general, when a tree-based model and a neural net are trained on the same training examples, their predictions should not differ substantially within the region covered by training examples. However, the two models tend to extrapolate very differently. Specifically, let $\boldsymbol{f}^{NET}$ denote the neural net and $\boldsymbol{f}^{TREE}$ denote a tree-based model, both trained with the same training examples. We know $\boldsymbol{m}$ is atypical if $\boldsymbol{f}^{NET}(\boldsymbol{m})$ and $\boldsymbol{f}^{TREE}(\boldsymbol{m})$ differ substantially. In this case, we display a warning to users.
It is worth mentioning that $\boldsymbol{f}^{NET}$ has a lower validation loss than $\boldsymbol{f}^{TREE}$. This result suggests that, within the region covered by training examples, the neural net outperforms the tree-based model. Therefore, in applications that are not deemed ill-suited, we use the estimates from the neural net instead of the tree-based model.