### v4

# largesymbols"0D largesymbols"0D

Abstract
We develop pre-trained estimators for structural econometric models. The estimator uses a neural net to recognize the structural model’s parameter from data patterns. Once trained, the estimator can be shared and applied to different datasets at negligible cost and effort. Under sufficient training, the estimator converges to the Bayesian posterior given the data patterns. As an illustration, we construct a pretrained estimator for a sequential search model (available at pnnehome.github.io). Estimation takes only seconds and achieves high accuracy on 12 real datasets. More broadly, pretrained estimators can make structural models much easier to use and more accessible.
1 Introduction
Pre-trained machine learning models, ranging from popular image recognition models to large language models, are achieving significant success (e.g., AlexNet, GPT). These pretrained models allow end users to greatly reduce the costs of pattern-recognition tasks. Take image recognition for example: instead of collecting images, labeling them, and training a neural net from scratch, users can directly apply a pretrained neural net to recognize objects in their images.
This paper explores the use of pretraining in structural estimation. We develop pretrained estimators for structural econometric models. The bulk of computational cost and researcher effort occur only once during the estimator’s construction. Subsequently, the estimator can be shared and applied to different datasets with negligible cost and effort. More specifically, a pretrained estimator assumes a given structural model. We use the structural model to simulate datasets, which are then used to train a neural net to recognize the structural model’s parameter from a set of data patterns (similar to how neural nets recognize objects from images). With sufficient training, the estimator converges to the Bayesian posterior of the parameter given the data patterns. We refer to the estimator as the pretrained neural net estimator (hereafter, pretrained NNE).
Pretrained NNE makes structural models easier to use and more accessible. In practice, estimating a structural model is often technically complex and time-consuming. Yet, the value of structural models – such as uncovering micro-founded preferences and enabling counterfactuals – is realized only after estimation. A pretrained NNE can be built once and then shared, allowing its users to easily and quickly estimate a structural model across applications. For users, the cost and effort of estimation become comparable to those in reduced-form regressions. Further, pretrained estimators can broaden the scope of applications of structural models. For example, one may apply pretrained estimators continuously to real-time data, enabling the integration of structural models into real-time algorithms (e.g., dynamic pricing, recommenders, bandit experiments). Another benefit of pretrained NNE is that it helps preserve privacy in estimation because it relies only on aggregate data patterns as input.
As an illustration, this paper focuses on sequential search models. These models characterize consumer search and purchase choices, and are finding growing application especially with online shopping data. Search models are estimated by simulated maximum likelihood (SMLE), but estimation is challenging. The vast number of search and purchase combinations makes it infeasible to directly simulate likelihood. To this end, the literature has introduced several workarounds over time, such as likelihood smoothing or specialized simulation algorithms (see Ursu et al. 2023). Even then, a single estimation run is expensive, let alone the multiple runs needed to tune optimizers, try starting points, tune smoothing factors, tune simulation algorithms, etc. Importantly, these tunings have to be performed separately for each application. More broadly, similar estimation challenges exist in many other structural models.
Pretrained NNE offers an alternative way to estimate a structural model. It does so by learning a direct mapping from data patterns to the structural model’s parameter. Conceptually, we can think of a pretrained NNE as follows. Research papers using structural estimation often start with reduced-form evidence before estimating the structural model. Although reduced-form results do not directly translate into an estimate of the structural model’s parameter, they offer useful clues. For example, if customers are more likely to search lower-price items, we expect the price coefficient in the search model to be more negative. If customers make more searches over weekends, we know weekend customers have either a low search cost or a low outside utility. More generally, we can imagine that with a rich set of such reduced-form patterns, an expert familiar with the search model can guesstimate the search model parameter value. The present paper’s exercise is to leverage machine learning to create such an “expert.”
Figure 1: An Illustration of NNE
We describe a standard sequential search model in Section 2. Then, in Section 3 we describe how to pretrain an NNE for this search model. At a high level, the training uses a large number of datasets simulated from the search model. Let $\boldsymbol{\theta}$ denote the search model parameter. For each draw of $\boldsymbol{\theta}$ from a prior, we can use the search model to simulate a dataset $\cal D$}. Use $\boldsymbol{m}$ to collect some data patterns of $\cal D$} (e.g., summary statistics, regression coefficients). Repeating this process for $L$ draws of $\boldsymbol{\theta}$ gives us $\{\boldsymbol{\theta}^{(\ell)},\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$ (we currently use $L\simeq$ 1 million). These $L$ pairs serve as the training examples for a neural net to learn a mapping from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. The trained neural net becomes the aforementioned “expert” capable of estimating the search model. Figure 1 offers an illustration of this neural net.
From a theoretical perspective, the estimate of NNE reflects a Bayesian posterior. Under mild conditions, as $L\rightarrow\infty$, the pretrained NNE converges to the posterior mean of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. This mean is known as a limited-information posterior mean, and it is one’s best estimate for the structural model’s parameter after seeing the data patterns.
We provide our pretrained NNE at: pnnehome.github.io. It can be applied off-the-shelf to estimate the search model: users only need to supply data, no further tuning is required, and the computation typically takes seconds on a laptop. Specifically, users supply a dataset $\cal D$} consisting of two components: (i) $\boldsymbol{y}$, which includes search and purchase choices, and (ii) $\boldsymbol{x}$, which includes product attributes that affect utility, consumer attributes that affect outside utility, and advertising attributes that affect search cost. The current version of our pretrained NNE can accommodate datasets with $n\geq 1000$ consumers, 2 to 8 product attributes, 0 to 5 consumer attributes, and 0 to 2 advertising attributes. These ranges are configurable before pretraining.
Section 4 evaluates the pretrained NNE using 12 real datasets. We conduct two sets of tests. The first set consists of Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices under a “true” $\boldsymbol{\theta}$. Pretrained NNE is faster than SMLE by 3-4 orders of magnitude. Further, it recovers $\boldsymbol{\theta}$ with accuracy equal to or higher than SMLE. The result holds even when compared against the SMLE that uses a specialized simulation algorithm for search models that takes considerable effort to implement (Jiang et al. 2021). The second set of tests estimates the search model entirely from real data. Because the true $\boldsymbol{\theta}$ is unknown, we compare the pretrained NNE’s estimates with SMLE’s estimates. Their estimates turn out to be similar. The pretrained NNE is again faster by 3-4 orders of magnitude.
An important detail in pretraining NNE is to specify the data patterns $\boldsymbol{m}$ — a representation of the input $\cal D$}. In computer vision or language processing, representations of inputs (images or texts) typically must be learned via training. By contrast, in structural models, economic theory and intuition often provide sufficient guidance for us to specify an $\boldsymbol{m}$ that can identify $\boldsymbol{\theta}$. This greatly reduces the training complexity. To optimize estimation efficiency, alternative specifications of $\boldsymbol{m}$ can be compared using the neural net’s accuracy on validation examples (i.e., a held-out subset of training examples). Importantly, this effort to select $\boldsymbol{m}$ arises only in pretraining and is not borne by the users of a pretrained NNE. For our search model, we choose $\boldsymbol{m}$ to include the coefficients from a set of regressions that describe consumer search and purchase behaviors. [^1] In Wei and Jiang (2024), $\boldsymbol{m}$ includes only data moments. Expanding $\boldsymbol{m}$ to include reduced-form regression coefficients leads to a substantial improvement in estimation accuracy for the search model.
This paper extends the standard (non-pretrained) NNE in Wei and Jiang (2024). In a standard NNE, all training examples assume the same observed attributes $\boldsymbol{x}$ from a real dataset. Thus, the trained net is applicable only to that dataset. In contrast, a pretrained NNE is applicable across different datasets. To achieve this, in each training example we draw $\boldsymbol{x}$ from a distribution (and then use the search model to simulate $\boldsymbol{y}$ given the drawn $\boldsymbol{x}$). This distribution serves the role of a prior for $\boldsymbol{x}$. We specify it to reflect attributes common in real applications, covering diverse scenarios that vary in the number of attributes, marginal distributions, correlations, and within- versus cross-consumer variation. We use a large $L$ to cover these scenarios well. It is important to note that an NNE (standard or pretrained) still assumes a specific structural model. Therefore, pretrained NNE is more suited for established structural models that are poised for broader applications, whereas standard NNE is more suited for newly proposed or ad hoc structural models.
Finally, we emphasize that this paper’s goal is to explore the feasibility of pretrained estimators for structural models, rather than to develop a full-fledged tool for estimating search models. For the latter task, there remains room for improvement, such as allowing richer unobserved consumer heterogeneity and incorporating instruments to address possible endogeneity. These extensions are non-trivial and we discuss them in Section 5.
1.1 Literature
This paper is related to several streams of research. First, this paper joins a recent literature that seeks to leverage machine learning for econometric estimation. In particular, the literature has studied the use of flexible machine learning models to better capture individual-level heterogeneity (e.g., Wager and Athey 2018, Athey et al. 2019, Farrell et al. 2021b), to better incorporate high-dimensional nuisance parameters (e.g., Chernozhukov et al. 2018), and to construct instrumental variables (e.g., Hartford et al. 2017, Lewis and Syrgkanis 2018, Singh et al. 2020). A main theme of this literature so far is to use machine learning to flexibly describe the relations between variables in data. The current paper differs in that the relations between data variables are described by an existing econometric model (e.g., sequential search), and we leverage machine learning to estimate the parameter of this econometric model. Recent works sharing this theme include Wei and Jiang (2024) and Kaji et al. (2023). In particular, Kaji et al. (2023) borrow ideas from adversarial learning, and propose training a discriminator to construct the objective function for estimating the econometric model. This approach is conceptually different from ours.
Second, we build on the successful idea of pretraining in machine learning practice. Pretraining involves training a machine learning model to learn general capabilities which can then be applied to various specific tasks. In natural language processing, pretrained models like BERT and GPT demonstrate remarkable capabilities in many language tasks (Devlin et al. 2018, Brown et al. 2020). In computer vision, pretrained neural nets such as AlexNet and ResNet have been applied widely (Krizhevsky et al. 2012, He et al. 2016). A review can be found in Han et al. (2021). Pretrained models have considerably improved the productivity of researchers and practitioners. The current paper aims to bring the benefits of pretraining to structural estimation in economics and marketing. Pretrained estimators are “Train Once, Deploy Anywhere.” Their availability will make structural estimation more accessible to researchers and practitioners.
Third, this paper adds to the literature that seeks to enable model estimation while preserving data privacy. Data privacy is becoming increasingly relevant to marketers. Various methods have been proposed to preserve privacy in estimation, such as noise infusion (Abowd et al. 2012, Avella-Medina 2021), generating synthetic data (Schneider et al. 2018, Anand and Lee 2023), and estimation using aggregate data (e.g., Musalem et al. 2009). The present paper advances privacy-preserving estimation using aggregate data. Pretrained NNE requires only reduced-form data patterns to estimate a structural model. Sharing reduced-form data patterns, which are aggregated across individuals, poses far less privacy concern for data owners than sharing individual-level data.
2 Consumer search model
In this section, we describe a standard (sequential) search model and the typical data used to estimate it. The next section will present the pretrained NNE for this search model.
2.1 The model
Sequential search models describe how consumers search and choose from a set of available options. One example where these models are increasingly applied is online browsing and shopping. In this context, consumer is given a webpage with a list of products; she decides which products to search (i.e., click) and which product to buy. A survey of sequential search models can be found in Ursu et al. (2023).
The current paper focuses on a fairly standard search model. Formally, there are $n$ consumers (or search sessions) and $J$ products per consumer. Let vector $\boldsymbol{x}_{ij}^{(prod)}$ collect the attributes of product $j$ for consumer $i$. The utility of consumer $i$ from product $j$ is:
$u_{ij}=\boldsymbol{\beta}^{\prime}\boldsymbol{x}_{ij}^{(prod)}+\mu_{ij}+\varepsilon_{ij}.$
There are two shocks, $\mu_{ij}$ and $\varepsilon_{ij}$. The consumer observes $\mu_{ij}$ pre-search but observes $\varepsilon_{ij}$ only post-search. Both shocks are distributed as $N(0,1)$. While normalizing one shock is without loss of generality, normalizing both shocks is not. Nevertheless, many applications have normalized both shocks because it can be empirically difficult to identify the variances of both shocks. We follow this practice. See Ursu et al. (2023) for more discussion.
Let vector $\boldsymbol{x}_{i}^{(cons)}$ collect the attributes of consumer $i$ (e.g., demographics, purchase history). These attributes affect the utility of an outside option as follows. The shock $\varepsilon_{i0}$ is known pre-search and follows $N(0,1)$.
$u_{i0}=\eta_{0}+\boldsymbol{\eta}^{\prime}\boldsymbol{x}_{i}^{(cons)}+\varepsilon_{i0}.$
Each consumer has one free search. Afterwards, a search cost is incurred for each additional search. We allow for some product-level attributes that affect search cost but not utility, collected in $\boldsymbol{x}_{ij}^{(ads)}$. We will refer to them as advertising or marketing attributes. One example of such attributes is the product rankings on the webpage. The search cost $c_{ij}$ is specified as
$\log(c_{ij})=\alpha_{0}+\boldsymbol{\alpha}^{\prime}\boldsymbol{x}_{ij}^{(ads)}.$
We use $d^{(prod)}$, $d^{(ads)}$, and $d^{(cons)}$ to denote the numbers of attributes in $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ , and $\boldsymbol{x}_{i}^{(cons)}$, respectively. Let $d\equiv d^{(prod)}+d^{(ads)}+d^{(cons)}$ denote the number of all attributes. Let $\boldsymbol{\theta}=(\boldsymbol{\beta}^{\prime},\boldsymbol{\eta}^{\prime},\boldsymbol{\alpha}^{\prime},\eta_{0},\alpha_{0})^{\prime}$ denote our search model’s parameter. The dimension of $\boldsymbol{\theta}$ is $d+2$.
The search model characterizes: (i) which products the consumer will search, (ii) which products the consumer will buy, and (iii) the order in which the consumer searches the products. We shall not delve into deriving the optimal choices of the consumer. It suffices to note that these choices are typically computed using the optimal decision rule in Weitzman (1979).
It is useful to note that the search model can also be applied to contexts outside online browsing and purchase. For example, geo-location data are increasingly available. They allow one to study consumers’ search behaviors in offline shopping. Shop visits and subsequent purchases can be modeled by a sequential search model. An interesting aspect is that the distances from shops to elevators or entrances may affect search costs (and thus be considered as a marketing attribute).
2.2 The data
Next, we describe the typical data used to estimate the search model. In doing so, we also make clear the range of data to which our pretrained NNE is applicable.
Continuing with the notation so far, we use $i$ to index consumers and $j$ to index products. Let $\boldsymbol{x}_{ij}^{(prod)}$ collect product attributes, $\boldsymbol{x}_{ij}^{(ads)}$ collect advertising attributes, and $\boldsymbol{x}_{i}^{(cons)}$ collect consumer attributes. A dataset $\cal D$} can be written as
${\cal D}$ $=\{\boldsymbol{x},\;\boldsymbol{y}\},$ (1)
$\text{where}:$ $\quad\boldsymbol{x}=\{\{\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)}\}_{j=1}^{J},\>\boldsymbol{x}_{i}^{(cons)}\}_{i=1}^{n},$
$\quad\boldsymbol{y}=\{\{\boldsymbol{y}_{ij}\}_{j=1}^{J}\}_{i=1}^{n}.$
Here, vector $\boldsymbol{y}_{ij}$ collects two outcome variables:
  • $y_{ij}^{(search)}$ is a dummy indicating whether the option $j$ is searched;
  • $y_{ij}^{(buy)}$ is a dummy indicating whether $j$ is bought.
In some applications, the order of searches is also observable, so $\boldsymbol{y}_{ij}$ can have a third element $y_{ij}^{(order)}$. However, because the search order is not always observable, in this paper we choose not to make use of search order information (so $y_{ij}^{(order)}$ is not required). Conceptually, it is straightforward to pretrain a separate NNE that uses the search order information.
Our pretrained NNE assumes that $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ and $\boldsymbol{x}_{i}^{(cons)}$ are all standardized to have zero means and unit variances. The means and variances here are computed across $i$ and $j$, e.g., the mean for $\boldsymbol{x}_{ij}^{(prod)}$ is $\frac{1}{nJ}\sum_{i=1}^{n}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$. If the attributes are not standardized, we standardize them before using the pretrained NNE to estimate $\boldsymbol{\theta}$. We then rescale the estimate of $\boldsymbol{\theta}$ to make it consistent with respect to the original, non-standardized attributes.
Table 1: Specified Ranges for Dataset Size
$d^{(prod)}$ $d^{(ads)}$ $d^{(cons)}$ $J$ $n$
[…]
max 8 2 5 35  $\begin{array}[]{c}\infty\;\text{ (application)}\\ 2e4\text{ (pretraining)}\end{array}$
The size of $\cal D$} is determined by the numbers of attributes ($d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$), number of products per consumer ($J$), and number of consumers ($n$). These numbers vary across applications. Our current version of pretrained NNE can be applied to datasets of sizes within the ranges specified in Table 1. These ranges are configurable before pretraining.
A note about Table 1 is that no cap is imposed on $n$ for applying the pretrained NNE. But a cap on $n$ is necessary in pretraining, because we cannot simulate data of infinite size. This cap is currently set at $2e4$. The pretrained NNE is not directly applicable to data larger than this cap. However, in practice, we can split a very large dataset and apply the pretrained NNE to each split. We then take the average of the estimates across the splits.
It is difficult to estimate $\boldsymbol{\theta}$ if the data contain too few searches or purchases. Therefore, we do not pretrain NNE to accommodate such data. Specifically, we look at:
  • Buy rate: fraction of consumers that bought some product (not outside good).
  • Search rate: fraction of consumers that did at least one non-free search.
We require that the buy rate is at least 0.5% and the search rate is at least 1%. Again, these thresholds are configurable prior to pretraining. During pretraining, simulated datasets where search rates or buy rates are lower than the thresholds are dropped.
As we will describe later, the pretraining incorporates non-normally distributed attributes. Nevertheless, the pretrained NNE may still perform poorly if certain attributes are too heavily skewed or have too many outliers. Therefore, we recommend using the usual practices to prepare data as in any empirical analysis. These include trimming, winsorizing, transforming variables, and dropping any attribute that causes (almost) collinearity.
3 Pretrained NNE
This section describes the construction of pretrained NNE. Although we aim to make the description self-contained, more emphasis is given to where the pretrained NNE differs from the standard NNE (Wei and Jiang 2024).
3.1 Overview
Let $\cal D$} denote a generic dataset as defined in expression (1). Let $\boldsymbol{\theta}$ be the parameter vector of the search model specified in Section 2.1. The task of NNE is to map $\cal D$} to an estimate of $\boldsymbol{\theta}$. It does so through two functions, $\boldsymbol{f}$ and $\boldsymbol{h}$.
$\widehat{\boldsymbol{\theta}}=\boldsymbol{f}\big[\>\underbrace{\boldsymbol{h}({\cal D})}_{\boldsymbol{m}}\>\big].$
  • The first function $\boldsymbol{h}$ calculates a set of data patterns of $\cal D$}, collected in vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$. This function is specified by us (rather than trained via examples).
  • The second function $\boldsymbol{f}$ is a neural net that maps from the data patterns to an estimate of $\boldsymbol{\theta}$. We write $\widehat{\boldsymbol{\theta}}=\boldsymbol{f}(\boldsymbol{m})$.
Conceptually, one could forgo specifying $\boldsymbol{h}$ and instead train a neural net mapping directly from $\cal D$} to $\boldsymbol{\theta}$. But the size of $\cal D$} is much larger than $\boldsymbol{m}$. In addition, the mapping from $\cal D$} to $\boldsymbol{\theta}$ tends to be much more complex than from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. So the specification of function $\boldsymbol{h}$ substantially reduces the complexity of the learning task. Further, the use of aggregate data patterns also allows privacy-preserving estimation, as discussed in Section 1.1.
In general, the vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ may include any data pattern of $\cal D$}. Ideally, we want to include data patterns that are most informative about $\boldsymbol{\theta}$ and yet inexpensive to compute. The foremost candidates are various data moments (hence the notation $\boldsymbol{m}$), such as means, variances, correlations. Further candidates are coefficients of reduced-form regressions. The idea of using reduced-form coefficients as “moments” originated from indirect inference (Gourieroux et al. 1993). Including reduced-form coefficients in $\boldsymbol{m}$ turns out to substantially improve the estimation accuracy of NNE, and is an important extension to Wei and Jiang (2024).
The detailed specification of $\boldsymbol{h}$ in our pretrained NNE is given in the appendix. Below, we give a summary.
  • First included in $\boldsymbol{m}$ are data size and summary statistics: $d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$, $n$, $J$, buy rate, search rate, the average number of searches per consumer.
  • To capture search decisions, we include the coefficients from a logit regression that predicts $y_{ij}^{(search)}$. Similarly, to capture purchase decisions, we include the coefficients from a multinomial logit regression that predicts $y_{ij}^{(buy)}$ conditional on search, i.e., $y_{ij}^{(search)}=1$. These regressions are at the product level.
  • We include the coefficients from a set of regressions that predict outcomes at the consumer level (e.g., whether consumer $i$ bought anything, whether $i$ made any non-free search).
Next, we turn to the function $\boldsymbol{f}$, which plays a central role in the pretrained NNE. The function is a neural net trained using simulated datasets. The training is conducted as follows.
  • We use the search model to simulate $L$ datasets, ${\cal D}^{(1)}$, ${\cal D}^{(2)}$, …, ${\cal D}^{(L)}$, each under a different parameter vector, $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$.
  • For each $\ell=1,2,...,L$, we compute the data patterns $\boldsymbol{m}^{(\ell)}=\boldsymbol{h}({\cal D}^{(\ell)})$.
  • Each pair $\{\boldsymbol{m}^{(\ell)},\boldsymbol{\theta}^{(\ell)}\}$ is a training example. We use $\{\boldsymbol{m}^{(\ell)},\boldsymbol{\theta}^{(\ell)}\}_{\ell=1}^{L}$ to train a neural net to predict $\boldsymbol{\theta}$ from $\boldsymbol{m}$.
We have left out three details in this training process: (i) the configuration of the neural net, (ii) the distribution from which we draw $\boldsymbol{\theta}^{(\ell)}$, and (iii) the distribution from which we draw the product, consumer, and advertising attributes in ${\cal D}^{(\ell)}$. These details are discussed in Section 3.3. Generally speaking, the distributions in (ii) and (iii) should be broad enough for NNE to accommodate a variety of real datasets in applications, but not so broad that NNE wastes resources learning scenarios that are unlikely in reality. A good balance makes the training more effective for a given $L$. Meanwhile, it is important for us to point out the pretrained NNE’s estimates in applications (Section 4) are not sensitive to the specifications of (i)-(iii).
The description so far has focused on point estimates. For the standard error of $\widehat{\boldsymbol{\theta}}$, we recommend bootstrapping. This boils down to applying the pretrained NNE to resampled copies of $\cal D$}. Because applying the pretrained NNE to each copy of $\cal D$} incurs very little cost, bootstrapping standard errors is not computationally expensive at all. [^2] The bootstrapping here omits the uncertainty in $\boldsymbol{f}$ resulting from the neural net training. However, given the large number of training examples ($L\simeq 1$ million), this uncertainty tends to be small relative to the sampling uncertainty of $\cal D$}.
Finally, a practical issue is how NNE behaves when the search model is misspecified for the dataset in a real application. NNE takes data patterns as input and outputs $\widehat{\boldsymbol{\theta}}$. This output is given so that, when the search model is simulated at $\widehat{\boldsymbol{\theta}}$, the resulting data patterns are close to the input data patterns. Importantly, this behavior does not require the input data patterns to be generated by the search model. In this sense, NNE is well-behaved under misspecification. [^3] Also see Section 3.2 that shows the pretrained NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$, which tends to be well-behaved under misspecification provided that the observed $\boldsymbol{m}$ is within the model-implied support of $\boldsymbol{m}$. However, this behavior does require the input data patterns to have positive probability under the search model and priors. If not, the input will lie outside the range of inputs seen in training and the neural net has to extrapolate. We call such cases “ill-suited” for the pretrained NNE, and it is useful to detect them in practice. We provide a detection method in the appendix.
3.2 Connection to Bayesian estimation
The eventual test for a pretrained model is to see how it performs in applications. We do so for the pretrained NNE in Section 4 and find that it provides accurate and sensible estimates. That said, from a theoretical standpoint, it is important to understand the econometric meaning of the pretrained NNE’s estimate. Below, we will explain that pretrained NNE has a familiar Bayesian interpretation, in that it calculates the posterior mean of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. In words, it provides the best estimate of the structural model’s parameter $\boldsymbol{\theta}$ given the data patterns $\boldsymbol{m}$.
Continuing with the notation in (1), let ${\cal D}=\{\boldsymbol{x},\boldsymbol{y}\}$ denote a dataset. In particular, $\boldsymbol{x}$ and $\boldsymbol{y}$ collect the observations across all consumers. The vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ denotes a set of data patterns. We start with a result with the standard (non-pretrained) NNE, studied in Wei and Jiang (2024). The main difference between standard NNE and pretrained NNE is that the former is trained for a specific real dataset. The training examples do not simulate $\boldsymbol{x}$ but directly use the observed $\boldsymbol{x}$ from that real dataset.
The standard NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ as $L\rightarrow\infty$ (see Wei and Jiang 2024). This result does not require large $n$ and holds for any value of $n$. Also note that we make it explicit that the expectation conditions on a specific $\boldsymbol{x}$. [^4] The notation of Wei and Jiang (2024) omits the conditioning on $\boldsymbol{x}$ because the paper assumes an underlying probability space that is already conditional on the observed $\boldsymbol{x}$. The $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ is known as a limited-information Bayesian posterior mean (e.g., Kim 2002), where “limited-information” indicates that the posterior relies on less information than the full data. To compare, the full-information posterior is $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{y},\boldsymbol{x})$. Ideally, we would like to know the full-information posterior, but when computing it is difficult, the limited-information posterior provides a practical alternative.
We now turn to the pretrained NNE. In this case, the training examples do not assume a specific $\boldsymbol{x}$, but instead draw $\boldsymbol{x}$ from a distribution. Consequently, the estimate by the pretrained NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ as $L\rightarrow\infty$, where the expectation no longer conditions on $\boldsymbol{x}$.
We provide a formal statement for this convergence result. To be precise, we consider the joint probability distribution between $\boldsymbol{\theta}$ and $\boldsymbol{m}$, $P(\boldsymbol{\theta},\boldsymbol{m})$. This joint distribution is implied by the search model, our specification of $\boldsymbol{m}$, and the distributions from which we draw $\boldsymbol{\theta}$ and $\boldsymbol{x}$ in pretraining. Specifically, in terms of probability densities,
$p(\boldsymbol{\theta},\boldsymbol{m})=\int p(\boldsymbol{m}|\boldsymbol{\theta},\boldsymbol{x})p(\boldsymbol{\theta})p(\boldsymbol{x})\>d\boldsymbol{x}.$ (2)
Here, $p(\boldsymbol{m}|\boldsymbol{\theta},\boldsymbol{x})$ is the likelihood of observing data patterns $\boldsymbol{m}$ under the search model. Terms $p(\boldsymbol{\theta})$ and $p(\boldsymbol{x})$ are densities from which we draw $\boldsymbol{\theta}$ and $\boldsymbol{x}$ in pretraining, and $p(\boldsymbol{\theta},\boldsymbol{x})=p(\boldsymbol{\theta})p(\boldsymbol{x})$ as we will draw them independently. With the probability density in equation (2), we have
$\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})=\frac{\int\boldsymbol{\theta}p(\boldsymbol{\theta},\boldsymbol{m})d\boldsymbol{\theta}}{\int p(\boldsymbol{\theta},\boldsymbol{m})d\boldsymbol{\theta}}.$
Each training example $\{\boldsymbol{\theta}^{(\ell)},\boldsymbol{m}^{(\ell)}\}$ is a draw from $P(\boldsymbol{\theta},\boldsymbol{m})$. We use $\boldsymbol{f}_{L}$ to denote the neural net trained with $L$ training examples. That is, $\boldsymbol{f}_{L}=\min_{\boldsymbol{\zeta}\in{\cal F}_{L}}\sum_{\ell=1}^{L}\lVert\boldsymbol{\zeta}(\boldsymbol{m}^{(\ell)})-\boldsymbol{\theta}^{(\ell)}\rVert^{2}$, where ${\cal F}_{L}$ is the same class of single-hidden-layer neural nets defined in Wei and Jiang (2024). The result below states that the function $\boldsymbol{f}_{L}$ converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ as a function of $\boldsymbol{m}$.
Proposition 1.
Suppose: (i) $\Theta$ is compact, (ii) $\boldsymbol{m}$ has a compact support, (iii) $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ is a continuous function of $\boldsymbol{m}$. We have $\int\lVert\boldsymbol{f}_{L}(\boldsymbol{m})-\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})\rVert^{2}dP(\boldsymbol{m})\rightarrow 0$ in probability as $L\rightarrow\infty$.
The continuity condition should hold in most structural models as it only requires the posterior mean not to change abruptly for a slight change in the observed data patterns. In comparison, the compactness conditions can be more restrictive. While the compactness of $\Theta$ is common in econometric settings, it technically precludes unbounded priors such as the normal distribution. In practice, we can trim the prior at extreme values to ensure a compact $\Theta$. Similarly, the support-compactness of $\boldsymbol{m}$ can be satisfied if we trim extreme data patterns in training. Nevertheless, in applications (Section 4), we find that the pretrained NNE’s estimates show very little difference whether trimming is used in pretraining. Finally, it is worth emphasizing that the conditions in Proposition 1 are sufficient but not necessary.
Given this convergence result, a natural question is how to interpret $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. Formally, it is another limited-information posterior mean. However, for a more intuitive perspective, consider a practical scenario where a researcher wants to apply the search model to a company’s data. For privacy reasons, the company provides the researcher with only aggregate data patterns $\boldsymbol{m}$, instead of the full data $\{\boldsymbol{x},\boldsymbol{y}\}$. In this case, the researcher’s best estimate for the search model parameter will be precisely $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. From equation (2), we see that for $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ to be well defined, the researcher should form a prior belief over $\boldsymbol{\theta}$, $p(\boldsymbol{\theta})$, and a prior belief over $\boldsymbol{x}$, $p(\boldsymbol{x})$. In the pretrained NNE, these priors are the distributions from which $\boldsymbol{\theta}$ and $\boldsymbol{x}$ are drawn during pretraining (details in Section 3.3).
3.3 Pretraining details
Section 3.1 gave an overview of our pretrained NNE, leaving out several details. We provide these details below. Readers interested mainly in the applications may skip these details without much loss of continuity.
3.3.1 The neural net
A neural net accepts fixed-length input and gives fixed-length output. But the search model parameter $\boldsymbol{\theta}$ has a size of $d+2$, where $d$ is the total number of attributes and it varies across different datasets. We address this issue through padding. Specifically, recall that $d\equiv d^{(prod)}+d^{(ads)}+d^{(cons)}$. Use subscript “max” to denote the maximum allowed value. For example, Table 1 sets $d_{max}^{(prod)}=8$ and $d_{max}=8+2+5=15$. We configure the neural net to always output $d_{max}+2$ values. Out of these output values, only $d+2$ values are useful. The rest are the padded “idle” values. Using ellipses to represent idle values, we can write the output as:
$(\kern 5.0pt\alpha_{0}\kern 5.0pt,\kern 5.0pt\boldsymbol{\alpha}^{\prime}\kern 5.0pt,\underbrace{\kern 5.0pt\cdot\cdot\cdot\kern 5.0pt}_{d_{max}^{(ads)}-d^{(ads)}},\kern 5.0pt\eta_{0}\kern 5.0pt,\kern 5.0pt\boldsymbol{\eta}^{\prime}\kern 5.0pt,\underbrace{\kern 5.0pt\cdot\cdot\cdot\kern 5.0pt}_{d_{max}^{(cons)}-d^{(cons)}},\kern 5.0pt\boldsymbol{\beta}^{\prime}\kern 5.0pt,\underbrace{\kern 5.0pt\cdot\cdot\cdot\kern 5.0pt}_{d_{max}^{(prod)}-d^{(prod)}}).$
During training, the idle output values are ignored by the loss function. In applications, the idle output values are discarded.
A similar arrangement applies to the neural net’s input. The length of $\boldsymbol{m}$ varies by the numbers of attributes in the data. For our current settings in Table 1, the maximal length of $\boldsymbol{m}$ is 164. Any shorter $\boldsymbol{m}$ is padded with zeros to this maximal length before we feed it into the neural net. The exact detail of $\boldsymbol{m}$ is given in the appendix.
We currently use $L\simeq$ 1e6 training examples (after dropping “corner” datasets with very low search or buy rates; see Section 2.2). We hold out $5e4$ examples for validation. Increasing $L$ further only marginally reduces the validation loss. We also use the hold-out validation to choose the neural net architecture. While Proposition 1 is formally stated with single-hidden-layer neural nets, in practice we find that using a few more hidden layers can offer lower validation losses. We end up using three hidden layers with 512, 256, and 256 nodes. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the neural net architecture.
3.3.2 The parameter prior
When generating the training datasets, we draw $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$ from a distribution of $\boldsymbol{\theta}$. As explained in Section 3.2, this distribution of $\boldsymbol{\theta}$ effectively plays the role of a prior for $\boldsymbol{\theta}$. We want the prior to be sufficiently broad to cover possible values of $\boldsymbol{\theta}$ in real applications, but not so broad that many training datasets become unrealistic and thus wasteful. Below, we explain the prior used in our pretrained NNE. We have also tried several alternative priors. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the prior specification.
Unlike Bayesian estimation on a single dataset, a unique challenge in specifying our prior is that we need to accommodate datasets with varying dimensions ($d^{(prod)}$, $d^{(cons)}$, $d^{(ads)}$). Consider key statistics such as buy rate and search rate. In real applications, there is no inherent reason for us to expect these key statistics to vary systematically across datasets of different dimensions. Therefore, across training datasets, we would like the key statistics not to vary systematically by data dimensions, either.
This desire rules out common priors such as $\beta_{k}\sim N(0,1)$ for $k=1,...,d^{(prod)}$. With this prior, the buy rates would be systematically higher in training datasets with larger $d^{(prod)}$. To see why, note this prior implies $\lVert\boldsymbol{\beta}\rVert^{2}\sim\chi_{d^{(prod)}}^{2}$ which has an expectation equal to $d^{(prod)}$. Consider a simple case where $\boldsymbol{x}_{ij}^{(prod)}\sim N(\boldsymbol{0},\boldsymbol{I})$. Then, the variance of $\boldsymbol{\beta}^{\prime}\boldsymbol{x}_{ij}^{(prod)}$ for any given $\boldsymbol{\beta}$ equals $\lVert\boldsymbol{\beta}\rVert^{2}$, which in expectation increases with $d^{(prod)}$. A larger variance makes it more likely for at least one product to be better than outside option, leading to a higher buy rate.
To this end, we specify a prior where $\lVert\boldsymbol{\beta}\rVert$ is invariant to $d^{(prod)}$:
$\boldsymbol{\beta}\sim\sigma\cdot\chi_{1}\cdot\boldsymbol{\nu},$ (3)
where $\sigma$ is a constant, $\chi_{1}$ is chi distributed, and $\boldsymbol{\nu}$ is a unit vector with a uniform distribution on the unit sphere. An implication of this prior is $\lVert\boldsymbol{\beta}\rVert^{2}\sim\sigma^{2}\cdot\chi_{1}^{2}$, which is invariant to $d^{(prod)}$. We use the same type of prior for $\boldsymbol{\eta}$ and $\boldsymbol{\alpha}$.
Intuitively, the prior (3) can also be understood as letting the overall impact of product attributes (as measured by $\lVert\boldsymbol{\beta}\rVert$) be invariant to $d^{(prod)}$. This means that when there are more product attributes, we expect that each individual attribute has a lower chance to have a large effect on utility. This belief seems to align with what we see in real applications, where typically only a few “key” product attributes have substantial effects on consumers’ search and purchase decisions.
The discussion so far has left out the priors for the intercepts, $\alpha_{0}$ and $\eta_{0}$, which directly affect search cost and outside utility. We use normal priors for these intercepts, with their means and variances specified to produce realistic search rates and buy rates in training datasets (neither too small nor too close to 1).
3.3.3 Attributes ($\boldsymbol{x}$)
The generation of training examples requires us to specify a distribution of $\boldsymbol{x}$. Specifically, to generate a training dataset ${\cal D}^{(\ell)}$, we first draw $\boldsymbol{x}^{(\ell)}$ from this distribution. Then, given $\boldsymbol{x}^{(\ell)}$ and $\boldsymbol{\theta}^{(\ell)}$, we use the search model to simulate $\boldsymbol{y}^{(\ell)}$, and let ${\cal D}^{(\ell)}=\{\boldsymbol{x}^{(\ell)},\boldsymbol{y}^{(\ell)}\}$. Conceptually, the distribution of $\boldsymbol{x}$ should be broad enough to cover common values of $\boldsymbol{x}$ in real applications, but not so broad that many training datasets become unrealistic and thus wasteful. Below, we describe the distribution of $\boldsymbol{x}$ currently used by us. It is important to note that the pretrained NNE’s estimates in applications (Section 4) are not sensitive to the specification of this distribution.
First, we draw the size of $\boldsymbol{x}$. The size is determined by $d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$, $J$, and $n$. These numbers are picked at random from their respective ranges to which we would like the pretrained NNE to be applicable. Our current choices of these ranges are given in Table 1 (in Section 2.2). In particular, the upper bound of $n$ in pretraining is currently set at 2e4. However, this cap does not restrict the applications of pretrained NNE, because we can always split a larger dataset and average the estimates across the splits (also discussed in Section 2.2).
Next, we draw the attributes in $\boldsymbol{x}$. In this regard, we incorporate several features of $\boldsymbol{x}$ commonly seen in real applications. The first feature is that many attributes show substantial consumer-level variations. For example, some consumers see prices systematically higher than other consumers. Therefore, we draw $\boldsymbol{x}$ with two components: a within-consumer component plus a cross-consumer component (the latter is also known as consumer fixed effects). The within-consumer component is ignored for $\boldsymbol{x}_{i}^{(cons)}$, which by definition varies only cross consumers.
The second feature is that individual attributes often follow non-normal distributions. To this end, we draw each attribute initially as normal and then transform it into one of the following non-normal distributions: (i) a dummy, (ii) a scale from 1 to 5, and (iii) a bell-shaped but skewed continuous distribution. In principle, more types of distributions could be included. However, when we do so, the pretrained NNE’s estimates in applications (Section 4) remain mostly unchanged. So these three types of distributions seem sufficient.
The third feature is that attributes are often correlated. To this end, we let the initial attribute draws follow multivariate normal distributions with randomly drawn correlation matrices. The correlations persist through subsequent transformation into non-normal distributions. For example, if two attributes are positively correlated when first drawn as normal, they will remain positively correlated when one of them is transformed into a dummy.
4 Applications
We evaluate the pretrained NNE’s performance with 12 real datasets. First, we briefly describe each dataset. Then, we conduct Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices. Finally, we estimate the search model on the real data entirely.
Before proceeding, we briefly describe the existing approaches to estimate search models, for readers unfamiliar with the literature. Standard SMLE with accept-reject simulations is infeasible for search models because it requires prohibitive computational costs. A common workaround in literature is to smooth the simulated likelihood function (e.g., Honka and Chintagunta 2016, Ursu 2018). However, estimates are sensitive to smoothing factors and tuning these factors requires costly Monte Carlo experiments. More recently, Jiang et al. (2021) and Chung et al. (2023) extend the Geweke-Hajivassiliou-Keane (GHK) simulation algorithm to the likelihood of search models. The GHK-based SMLE does not require smoothing, but is much more complex to implement. In this section, we implement both the smoothed SMLE and the GHK-based SMLE.
4.1 Datasets
Our 12 datasets come from four different sources (an anonymous online retailer, Expedia, JD, and Trivago). Table 2 provides summary statistics. We see that these datasets vary substantially in the number of search sessions, numbers of attributes, buy rate, and search rate. So they cover a fairly wide range of real-world scenarios.
For ease of presentation, the estimates reported in this section assume that each attribute is mean-zero with a standard deviation of 0.2. The 0.2 standard deviation (instead of 1) makes coefficients ($\boldsymbol{\beta}$, $\boldsymbol{\eta}$, $\boldsymbol{\alpha}$) larger and thus more visible in figures. For estimation, all attributes are standardized to unit variance before being fed into the pretrained NNE (see Section 2.2).
We briefly describe each dataset below.
Table 2: Summary Statistics of 12 Datasets
$n$ $d^{(prod)}$ $d^{(cons)}$ $d^{(ads)}$ buy rate search rate
eCommerce: laptop 10,000 6 0 0 5.9% 21.5%
eCommerce: vacuum 10,000 4 0 0 8.5% 19.0%
eCommerce: washer 10,000 5 0 0 7.8% 22.1%
Expedia: destination 1 1,258 6 2 1 11.5% 8.7%
Expedia: destination 2 897 6 2 1 3.8% 6.0%
Expedia: Cancun 2,135 6 0 1 1.3% 30.8%
Expedia: Manhattan 1,369 6 0 1 3.1% 28.7%
JD: desktop channel 5,896 4 2 0 16.4% 26.5%
JD: mobile channel 4,069 4 2 0 11.2% 12.2%
JD: app channel 10,000 4 2 0 8.0% 12.9%
Trivago: desktop channel 2,082 7 0 1 2.8% 40.0%
Trivago: mobile channel 2,726 7 0 1 2.5% 29.1%
Notes: $n$ is the number of consumers or search sessions. $d^{(prod)}$, $d^{(cons)}$, and $d^{(ads)}$ are the numbers of product attributes, consumer attributes, and advertising attributes, respectively. Buy rate is the percentage of consumers who made a purchase. Search rate is the percentage of consumers who made at least one non-free search.
(1) eCommerce: laptop
Our first data source consists of consumer searches and purchases in an anonymous large online store, provided by the marketing platform REES46. The data are available on Kaggle.com and include many product categories. [^5] https://www.kaggle.com/datasets/mkechinov/ecommerce-behavior-data-from-multi-category-store We focus on laptops in this dataset. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and five brand dummies: Asus, HP, Lenovo, MSI, and Xiaomi (the baseline brand is Acer). There are no consumer or advertising attributes. There are a very large number of search sessions. To ease the computational burden of SMLE, we use a subsample of 10,000 search sessions.
(2) eCommerce: vacuum
This dataset includes the vacuum category in the anonymous online store. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and three brand dummies: Philips, Samsung, and Vitek (the baseline brand is Bosch). We again take a subsample of 10,000 search sessions.
(3) eCommerce: washer
This dataset includes the washer category in the anonymous online store. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and four brand dummies: Beko, Candy, Indesit, Midea (the baseline brand is Atlant). We again take a subsample of 10,000 search sessions.
(4) Expedia: destination 1
The International Conference on Data Mining (ICDM 2013) organized a contest to predict consumer searches and bookings for hotels on Expedia. The data for the contest are publicly available on Kaggle.com. [^6] https://www.kaggle.com/competitions/expedia-personalized-sort/overview Several papers have used the data to study consumer search behaviors (Ursu 2018, Greminger 2022, Compiani et al. 2023). The data include multiple travel destinations. “Destination 1” refers to the largest destination. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include star rating, review score, a chain indicator, location score, price, and promotion. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The advertising attribute $\boldsymbol{x}_{ij}^{(ads)}$ includes the ranking of the hotel.
(5) Expedia: destination 2
This dataset consists of the search sessions for the second-largest destination in the Expedia data from ICDM 2013. There are only $n=897$ search sessions, slightly below the lower bound of $n$ set for the pretrained NNE (see Table 1). Nonetheless, we choose to include the dataset, because it allows us to test the pretrained NNE slightly outside its intended range of applications.
(6) Expedia: Cancun
The Analytics at Wharton Initiative also provides data on hotel searches and bookings on Expedia. [^7] https://analytics.wharton.upenn.edu/open-research-opportunities/predicting-optimizing-consumer-response-to-product-search-results/ The data are not openly available but can be accessed after a request with the Initiative. The data include four travel destinations (Cancun, Manhattan, Paris and Budapest). Cancun is the largest destination.
(7) Expedia: Manhattan
This dataset consists of the search sessions for the second-largest destination (Manhattan) in the Expedia data from Analytics at Wharton.
(8) JD: desktop channel
The 2020 MSOM Data Driven Research Challenge provided data on consumer experiences on JD.com. The data were made available to INFORMS members in 2019.888https://connect.informs.org/msom/events/datadriven2020 JD.com is China’s second largest business-to-consumer online retail platform. The data cover many aspects of consumer experiences on the platform, from website browsing to order fulfillment (see Shen et al. 2020 for details). We focus on consumer search and purchase activities. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price, a dummy for first-party sellers, and two unnamed attributes. There are no advertising attributes.
The JD.com data include three channels: desktop, mobile, and app. This dataset focuses on the desktop channel.
(9) JD: mobile channel
This dataset includes the mobile channel from the JD.com data.
(10) JD: app channel
This dataset includes the app channel from the JD.com data. There are a large number of search sessions. To ease the computational burden of SMLE, we use a subsample of 10,000 search sessions.
(11) Trivago: desktop channel
The ACM RecSys Challenge 2019 provided consumer browsing data on Trivago.com. [^9] https://recsys.acm.org/recsys19/challenge/ The data are not standard for search models because purchases are not very well defined in this context, as we will explain below. Nonetheless, we choose to include the data because doing so allows us to test the pretrained NNE slightly outside its intended context of applications.
Trivago is a hotel meta-search engine that aggregates information from hotels and travel websites (similar to how Google Flights aggregates flight information). On Trivago.com, consumers receive a list of hotels. They can click on any hotel to see more information. They can also click a “View Deal” button that directs them to an external website for possible booking. But what happens at external websites is not observed in the data. We treat a click for more information as a search and a “View Deal” click as a purchase. Sometimes a consumer clicks “View Deal” for multiple hotels, in which case we randomly pick one as the purchase. Cases where consumers click “View Deal” without first clicking for more information are dropped (about 5%).
The Trivago data include search sessions from both the desktop channel and mobile channel. This dataset focuses on the desktop channel.
(12) Trivago: mobile channel
This dataset consists of the search sessions from the mobile channel in the Trivago data.
4.2 Monte Carlo studies
Table 3: Estimates in Monte Carlo Study with eCommerce: Laptop
Truth Pretrained NNE SMLE (GHK) $R=50$ SMLE (smoothed) $R=50$ $\boldsymbol{\lambda}=(4,1)$ SMLE (smoothed) $R=50$ $\boldsymbol{\lambda}=(2,1)$ SMLE (smoothed) $R=50$ $\boldsymbol{\lambda}=(4,2)$
$\alpha_{0}$ -5.0 -5.009 (0.006) -5.779 (0.010) -4.696 (0.023) -3.347 (0.011) -2.767 (0.011)
$\eta_{0}$ 4.8 4.797 (0.002) 4.732 (0.004) 4.851 (0.009) 4.981 (0.005) 3.999 (0.006)
$\beta_{1}$ -0.6 -0.603 (0.002) -0.579 (0.003) -0.375 (0.005) -0.457 (0.005) -0.366 (0.005)
$\beta_{2}$ -0.5 -0.488 (0.002) -0.485 (0.003) -0.300 (0.005) -0.370 (0.005) -0.295 (0.005)
$\beta_{3}$ -0.3 -0.292 (0.002) -0.290 (0.003) -0.188 (0.005) -0.228 (0.006) -0.183 (0.005)
$\beta_{4}$ -1.0 -1.010 (0.003) -0.968 (0.003) -0.627 (0.005) -0.771 (0.005) -0.616 (0.005)
$\beta_{5}$ -0.4 -0.399 (0.003) -0.390 (0.003) -0.255 (0.005) -0.315 (0.006) -0.250 (0.005)
$\beta_{6}$ -0.5 -0.489 (0.003) -0.479 (0.003) -0.316 (0.006) -0.392 (0.006) -0.312 (0.005)
Notes: $R$ is the number of simulations per consumer in SMLE. $\boldsymbol{\lambda}$ is the smoothing factors in the smoothed SMLE. Results are averaged across 100 Monte Carlo repetitions. Numbers in parentheses are the standard errors for the averages.
Our first set of tests are Monte Carlo studies. Specifically, for each of the 12 real datasets, we: (i) keep the product, consumer, and advertising attributes, but (ii) replace the search and purchase choices with the choices simulated by the search model under a “true” $\boldsymbol{\theta}$. Then, we try to recover $\boldsymbol{\theta}$ from this dataset. Knowing the ground truth of $\boldsymbol{\theta}$ allows us to measure estimation accuracy. We set the true $\boldsymbol{\theta}$ close to its estimate from Section 4.3.
Figure 2: Monte Carlo Study with eCommerce: Laptop
Notes: Results are based on 100 Monte Carlo repetitions.
Below, we first closely examine the results from the eCommerce: laptop dataset. In Table 3, Column 1 shows the true value of $\boldsymbol{\theta}$, Column 2 shows the pretrained NNE’s estimate, and Columns 3 - 6 show the SMLE’s estimates. The $R$ denotes the number of simulations per consumer in SMLE and $R=50$ is typical in literature. For the smoothed SMLE, $\boldsymbol{\lambda}$ collects the smoothing factors for search choices and for purchase choices. [^10] Some applications use the same smoothing factor for search and purchase choices. But doing so here worsens RMSE substantially. Whether multiple smoothing factors are necessary is generally application-dependent. The table shows three settings of $\boldsymbol{\lambda}$ to illustrate the sensitivity of SMLE’s estimate to $\boldsymbol{\lambda}$. The optimal $\boldsymbol{\lambda}$ is generally application-dependent. We actually have conducted an expensive grid search of $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}=(4,1)$ gives the smallest root mean square error (RMSE) in this application.
The upper plot of Figure 2 reports the RMSEs of the different estimators. The pretrained NNE has the lowest RMSE, outperforming even the GHK-based SMLE that uses a very large $R=1000$. Although we generally expect MLE to be asymptotically efficient, the performance of SMLE is affected by simulation errors. In search models, simulating the likelihood is particularly difficult because of the large number of possible search and purchase combinations. Our result here suggests that simulating this likelihood remains challenging despite workarounds such as smoothing and GHK. The result also shows that pretrained NNE can achieve high estimation accuracy despite being off-the-shelf.
The lower plot of Figure 2 reports the computation time needed for a single estimation run. All the times are measured on the same hardware. Recall that this dataset is of moderate size (a subsample from a single product category; see Section 4.1). [^11] There are 624 categories in the original data. The laptop category has 41,918 search sessions. Our dataset is a subsample of 10,000 search sessions. Even so, the SMLE requires between 1.5 and 10 hours. The pretrained NNE takes about 1 second (so the orange bar is barely visible). This represents a time saving of 3 to 4 orders of magnitude. In addition, we note that implementing SMLE requires additional costs not reflected in Figure 2, such as tuning smoothing factors, tuning optimization routines, trying starting points. Each of these tasks requires multiple estimation runs. Moreover, the GHK takes considerable effort to code. While pretraining an NNE also has its own implementation costs, they arise only once during pretraining and are not borne by the subsequent users of the pretrained NNE.
Next, we expand the examination to all 12 datasets. Table 4 reports the results. The table omits the smoothed SMLE because, as we have seen, its estimates are sensitive to the smoothing factors $\boldsymbol{\lambda}$ (and the optimal $\boldsymbol{\lambda}$ is application-dependent). The main finding with Table 4 is qualitatively the same as what we have seen with the eCommerce: laptop dataset. The pretrained NNE achieves either comparable or better estimation accuracy at drastically lower computational costs.
Finally, it is worth noting that the 12 datasets here are not large (see Table 2). In fact, recall that each dataset focuses on only a subset (a single product category, travel destination, or sales channel) of its source, and four of the 12 datasets are subsampled to reduce the computation burden. Because computation time scales with data size, the absolute time savings on large datasets will be much greater than those indicated by Table 4.
Table 4: Monte Carlo Studies Across 12 Datasets
Pretrained NNE SMLE (GHK) $R=50$ SMLE (GHK) $R=1000$
RMSE
eCommerce: laptop 0.094 (0.003) 0.795 (0.010) 0.453 (0.009)
eCommerce: vacuum 0.095 (0.003) 0.352 (0.005) 0.176 (0.005)
eCommerce: washer 0.179 (0.009) 0.614 (0.007) 0.361 (0.008)
Expedia: destination 1 0.451 (0.012) 0.591 (0.022) 0.501 (0.015)
Expedia: destination 2 0.748 (0.023) 0.964 (0.035) 0.833 (0.028)
Expedia: Cancun 0.485 (0.021) 1.814 (0.055) 1.083 (0.045)
Expedia: Manhattan 0.500 (0.018) 1.449 (0.054) 0.855 (0.040)
JD: desktop channel 0.175 (0.005) 0.392 (0.008) 0.256 (0.008)
JD: mobile channel 0.234 (0.007) 0.375 (0.011) 0.273 (0.009)
JD: app channel 0.161 (0.006) 0.335 (0.006) 0.200 (0.008)
Trivago: desktop channel 0.423 (0.016) 2.462 (0.050) 1.521 (0.044)
Trivago: mobile channel 0.422 (0.021) 1.067 (0.037) 0.660 (0.030)
Compute time (mins)
eCommerce: laptop 0.02 132.72 630.06
eCommerce: vacuum 0.02 59.42 371.52
eCommerce: washer 0.02 126.48 552.85
Expedia: destination 1 0.01 20.35 126.20
Expedia: destination 2 0.01 17.53 94.03
Expedia: Cancun 0.02 27.13 151.97
Expedia: Manhattan 0.02 19.01 101.17
JD: desktop channel 0.02 50.43 306.65
JD: mobile channel 0.01 33.53 213.45
JD: app channel 0.02 85.27 495.20
Trivago: desktop channel 0.02 33.98 212.17
Trivago: mobile channel 0.01 46.50 281.87
Notes: $R$ is the number of simulations per consumer in SMLE. $R=50$ is typical in literature. Results are averaged across 100 Monte Carlo repetitions. Numbers in parentheses are the standard errors for the averages.
4.3 Real-data estimates
Figure 3: Real-Data Estimates Across 12 Datasets
eCommerce: laptop
eCommerce: vacuum
eCommerce: washer
Expedia: destination 1
Expedia: destination 2
Expedia: Cancun
Expedia: Manhattan
JD: desktop
JD: mobile
JD: app
Trivago: desktop
Trivago: mobile
Our second set of tests estimates the search model entirely from real data – that is, both real attributes and real consumer choices. We no longer have a ground truth for $\boldsymbol{\theta}$ and therefore cannot measure estimation accuracy directly. Instead, we assess whether the pretrained NNE’s estimates are sensible, by comparing them with the estimates from the GHK-based SMLE that uses the very large $R=1000$.
Figure 3 shows the estimates across the 12 datasets. Each plot corresponds to one dataset. The horizontal axis lists the individual parameters in $\boldsymbol{\theta}$. The orange bars show the pretrained NNE’s estimates, while the blue bars show the SMLE’s estimates. A 95% confidence interval is shown on each bar. Overall, we see that the estimates from the pretrained NNE and from the SMLE are closely matched.
We omit reporting computation times in Figure 3, because they are quantitatively similar to those reported in Table 4 for Monte Carlo studies. Across all the datasets, the pretrained NNE uses only a fraction of the computation time required by SMLE.
5 Conclusion
This paper explores the feasibility of pretraining an off-the-shelf machine learning model to estimate a given structural econometric model. We do not use the machine learning model to capture relationships in data. Instead, these relationships are prescribed by the structural model, and we use the machine learning model to recover this structural model’s parameter. Pretrained NNE enables users to estimate a structural model as easily as reduced-form regressions, while providing equal or better accuracy than existing estimation methods (at least in the context of the search model).
Structural estimation is very valuable in empirical research — it offers insights into micro-founded consumer preferences and enables counterfactual analyses. However, its wider use has been limited by the large amount of effort and time required to estimate structural models. Pretrained estimators remove these barriers, making structural models more accessible to both researchers and practitioners.
Further, the ease of computation offered by pretrained NNE may facilitate new applications. One notable area is to integrate structural models with real-time algorithms, such as dynamic pricing, ranking algorithms, recommender systems, and bandit experiments (e.g., Agrawal et al. 2019, Jain et al. 2024). Because structural models provide estimates of latent consumer preferences, these estimates can be fed into real-time algorithms to improve their decisions, e.g., what price to set or what items to recommend. In addition, counterfactual analyses can be used to evaluate and provide feedback on the algorithms. However, in such applications the structural models must be estimated quickly in real time. Pretrained NNE offers a way to achieve such fast estimation.
The intention of this paper is to explore the feasibility of pretrained estimators, rather than to develop a comprehensive estimator for search models. This leaves areas where the pretrained NNE can be further improved. We note two potential areas. The first area is to increase the richness of the structural model. The search model considered in this paper is a standard yet simple one. Future work may allow for attributes that simultaneously affect utility and search costs, or allow for richer unobserved heterogeneity in utility and search costs. These extensions will require us to find additional reduced-form patterns to help identify the more complex structural model.
A second area for improvement is to address potential endogeneity. For example, if researchers have an instrumental variable (IV) for price, how do we build a pretrained estimator that can leverage the instrument? At a high level, addressing this problem requires two tasks. The first task is to generate training datasets that include instruments. The second task is to pass relevant information regarding the instrument to the neural net. These tasks are not straightforward. However, incorporating IV in NNE will be a significant and valuable extension, because we frequently need to address endogeneity in structural estimation applications.
Appendix
Appendix A Data patterns
Section 3.1 summarized the data patterns $\boldsymbol{m}$ that we feed into the neural net. Below, we provide the exact specification of $\boldsymbol{m}$.
We first recall some notation that will be needed. Vectors $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$, and $\boldsymbol{x}_{i}^{(cons)}$ collect the product, advertising, and consumer attributes, respectively. It is assumed that all attributes are standardized. As before, let $d^{(prod)}$, $d^{(ads)}$, and $d^{(cons)}$ denote the numbers of attributes. These numbers are dataset-dependent. Let $d_{max}^{(prod)}$, $d_{max}^{(ads)}$, and $d_{max}^{(cons)}$ denote the maximum values that we currently set for the numbers of attributes (given in Table 1). These maximum values are not dataset-dependent.
We now define several consumer-level variables that will be used. Recall that $J$ is the number of products per consumer and $n$ is the number of consumers.
  • Let $\bar{\boldsymbol{x}}_{i}^{(prod)}=\frac{1}{J}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$ denote the average product attributes for consumer $i$.
  • Let $\bar{\boldsymbol{x}}_{i}^{(ads)}=\frac{1}{J}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(ads)}$ denote the average advertising attributes for consumer $i$.
  • Let $\widetilde{y}_{i}^{(search)}=\sum_{j=1}^{J}y_{ij}^{(search)}$ denote the number of searches made by consumer $i$.
  • Let $\widetilde{y}_{i}^{(nf)}=\mathbb{I}\big\{\widetilde{y}_{i}^{(search)}>1\big\}$ indicate whether consumer $i$ did any non-free search.
  • Let $\widetilde{y}_{i}^{(buy)}=\sum_{j=1}^{J}y_{ij}^{(buy)}$, which is a dummy indicating whether consumer $i$ made a purchase.
The elements of $\boldsymbol{m}$ for any given dataset $\cal D$} are:
  1. 1.
    The buy rate, search rate, and average number of searches per consumer.
  2. 2.
    The values of $J$ and $n$.
  3. 3.
    A dummy vector indicating the numbers of product, advertising, and consumer attributes, shown as follows. Note that the vector has a fixed length of $d_{max}\equiv d_{max}^{(prod)}+d_{max}^{(ads)}+d_{max}^{(cons)}$.
    $(\quad\underbrace{1,...,1,}_{d^{(prod)}}\quad\underbrace{0,...,0,}_{d_{max}^{(prod)}-d^{(prod)}}\quad\underbrace{1,...,1,}_{d^{(ads)}}\quad\underbrace{0,...,0,}_{d_{max}^{(ads)}-d^{(ads)}}\quad\underbrace{1,...,1,}_{d^{(cons)}}\quad\underbrace{0,...,0}_{d_{max}^{(cons)}-d^{(cons)}}).$ (4)
  4. 4.
    The coefficients for $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $y_{ij}^{(search)}$ on $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\allowbreak\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(prod)},\bar{\boldsymbol{x}}_{i}^{(ads)}\}$. These coefficients are padded with zeros and collected in a vector of fixed length $d_{max}+1$. The paddings enter this vector in the same fashion as in expression (4). This style of collecting regression coefficients applies to all the regressions below.
  5. 5.
    The coefficients in a multinomial logit regression of $y_{ij}^{(buy)}$ on $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)}\}$, conditional on $y_{ij}^{(search)}=1$. Intuitively, this regression describes how each consumer chooses among searched products.
  6. 6.
    The coefficients for $\{1,\;\boldsymbol{x}_{i}^{(cons)}\}$ in a linear regression of $\log(\widetilde{y}_{i}^{(search)})$ on $\{1,\;\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(ads)},\bar{\boldsymbol{x}}_{i}^{(prod)}\}$.
  7. 7.
    The coefficients for $\{1,\;\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $\widetilde{y}_{i}^{(nf)}$ on $\{1,\;\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(ads)},\bar{\boldsymbol{x}}_{i}^{(prod)}\}$.
  8. 8.
    The coefficients for $\{1,\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $\widetilde{y}_{i}^{(buy)}$ on $\{1,\;\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(ads)},\bar{\boldsymbol{x}}_{i}^{(prod)}\}$.
  9. 9.
    The means and standard deviations of the variables in all the regressions above, except for those with fixed values (e.g., the standard deviations of $\boldsymbol{x}_{ij}^{(prod)}$ always equal to 1).
All the regressions above are estimated using ridge regression that imposes a small penalty on the coefficients. Using ridge regression leads to a noticeable decrease in the validation loss of the pretrained NNE. The validation loss is further decreased when we estimate each regression with two different penalty levels, and include the coefficients from both penalty levels in $\boldsymbol{m}$. We currently use the penalty levels of 0.001 and 1e-6. Using three or more penalty levels yields little additional decrease in the validation loss.
Appendix B Detecting ill-suited applications
As discussed, we say a dataset is ill-suited if the associated data patterns $\boldsymbol{m}$ lies outside the range of data patterns seen in pretraining. In this case, the neural net has to extrapolate and its output may not be reliable. Because a pretrained NNE can be applied at low cost to many datasets, it is helpful to have a way to detect whether a particular dataset is ill-suited and to inform users when this is the case.
Specifically, we want to detect whether the $\boldsymbol{m}$ from a given real dataset is atypical (an outlier) relative to $\{\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$ from the training examples. We note that $\boldsymbol{m}$ has over a hundred dimensions, and the detection cannot be reduced to checking each dimension separately. Even if every individual element of $\boldsymbol{m}$ appears typical on its own, the vector $\boldsymbol{m}$ can still be atypical if it contradicts some relationship between elements that holds in the training examples.
We propose a practical method of detection by making use of a tree-based model (boosted tree or random forest). In general, when a tree-based model and a neural net are trained on the same training examples, their predictions should not differ substantially within the region covered by training examples. However, a tree-based model and a neural net tend to extrapolate very differently. So, let $\boldsymbol{f}^{NET}$ denote the neural net and $\boldsymbol{f}^{TREE}$ denote a tree-based model, both trained with the same training examples. We know $\boldsymbol{m}$ is atypical if $\boldsymbol{f}^{NET}(\boldsymbol{m})$ and $\boldsymbol{f}^{TREE}(\boldsymbol{m})$ differ substantially. In this case, we display a warning to users.
It is worth pointing out that $\boldsymbol{f}^{NET}$ has a lower validation loss than $\boldsymbol{f}^{TREE}$. In other words, within the region covered by training examples, the neural net outperforms the tree-based model. Therefore, in any application that is not deemed ill-suited, we use the output from the neural net instead of the tree-based model.
Appendix C Proof
Proposition 1 is proved by following Wei and Jiang (2024)’s proof (which builds on Chen 2007 and White 1989). The only difference is that Wei and Jiang (2024) work with a distribution of $(\boldsymbol{\theta},\boldsymbol{m})$ conditional on some observed $\boldsymbol{x}$, whereas Proposition 1 in the current paper considers the distribution of $(\boldsymbol{\theta},\boldsymbol{m})$ unconditional on $\boldsymbol{x}$. With this adjustment, the same argument in Wei and Jiang (2024) applies directly, so we omit repeating the details here.